### v2

# Hilbert schemes on blowing ups and the free Boson

Abstract.
For different cohomology theories (including the Hochschild homology, Hodge cohomology, and Chow groups), we identify the cohomology of moduli space of rank $1$ perverse coherent sheaves on the blow-up of a surface by Nakajima-Yoshioka [22] as the tensor product of cohomology of Hilbert schemes with a Fermionic Fock space. As the stable limit, we identify the cohomology of Hilbert schemes of blow-up of a surface as the tensor product of cohomology of Hilbert schemes on the surface with a Bosonic Fock space. The actions of the infinite-dimensional Clifford algebra and Heisenberg algebra are all given by geometric correspondences.
1. Introduction
1.1. Motivation and the Main Theorem
Let $S$ be a smooth projective surface over $\mathbb{C}$. The Poincaré polynomials of Hilbert schemes of points on $S$ were computed by the famous Göttsche’s formula:
(1.1) $\sum_{n=0}^{\infty}q^{n}P_{t}(S^{[n]})=\prod_{m=1}^{\infty}\frac{(1+t^{2m-1}q^{m})^{b_{1}(S)}(1+t^{2m+1}q^{m})^{b_{3}(S)}}{(1-t^{2m-2}q^{m})^{b_{0}(S)}(1-t^{2m}q^{m})^{b_{2}(S)}(1-t^{2m+2}q^{m})^{b_{4}(S)}},$
where $b_{i}(S)$ is the $i$-th Betti number.
  1. (1)
    By letting $t=1$, the formula above is the character of a super-Fock space. Nakajima [21] and Grojnowski [10] realized the cohomology group of the Hilbert scheme as the Fock space of a super-Heisenberg algebra through geometric correspondences.
    By replacing the cohomology with the Chow group, the Heisenberg (and moreover, Virasoro and $W_{1+\infty}$) algebra action was constructed by Maulik-Negut [19]. However, as the Chow group of a surface can be infinite-dimensional (like $K3$ surfaces), the Chow groups of all $S^{n}$ are invovled to decide the Chow group of Hilbert schemes, and the representation theory is also more complicated (for example, $sp_{2\infty}$ will naturally appear).
  2. (2)
    Let $\hat{S}$ be the blow-up of $S$ at a point $o$. By Göttsche’s formula, we have
    $\sum_{n=0}^{\infty}q^{n}P_{t}(\hat{S}^{[n]})=\prod_{m=1}^{\infty}\frac{1}{1-t^{2m}q^{m}}\sum_{n=0}^{\infty}q^{n}P_{t}(S^{[n]}),$
    where the ratio
    $\prod_{m=1}^{\infty}\frac{1}{1-t^{2m}q^{m}}$
    is the character of a Bosonic Fock space. Thus, it suggests we have a canonical isomorphism
    (1.2) $\bigoplus_{n=0}^{\infty}H^{*}(\hat{S}^{[n]},\mathbb{Q})\cong\bigoplus_{n=0}^{\infty}H^{*}(S^{[n]},\mathbb{Q})\otimes_{\mathbb{Q}}B,$
    where $B$ is the bosonic Fock space $\mathbb{Q}[y_{1},y_{2},\cdots]$. The formula (1.2) does not depend on the choice of the surface $S$ and the point $o$, and might also hold if we replace the cohomology group with Chow groups and Hochschild homology.
The purpose of this paper is to realize the expectation (1.2) by proving that:
Theorem 1.1 (Theorem 6.3).
Let $\mathbb{H}^{*}$ be the functor of Hodge cohomology groups, Hochschild homology groups, or the Chow groups. We have an identification
(1.3) $\bigoplus_{n=0}^{\infty}\mathbb{H}^{*}(\hat{S}^{[n]},\mathbb{Q})\cong\bigoplus_{n=0}^{\infty}\mathbb{H}^{*}(S^{[n]},\mathbb{Q})\otimes_{\mathbb{Q}}B$
where the Heisenberg algebra action on $B$ is induced by geometric correspondences.
Remark 1.2.
More details of our geometric correspondences will be given from Section 1.2 to Section 1.4. At the current stage, we do not know the precise relation between our Heisenberg algebra action with that of Nakajima [21] and Grojnowski [10] in cohomlogy groups or Maulik-Neguţ [19] in Chow groups.
1.2. Stable limits and perverse coherent sheaves
The right way to consider Theorem 1.1 is to regard the Bosonic Fock space $B$ and $\mathbb{H}^{*}(\hat{S}^{[n]},\mathbb{Q})$ as a stable limit:
  1. (1)
    We consider the infinite dimensional vector space $V$ and the Fermionic Fock space
    $V=\bigoplus_{i=1}^{\infty}\mathbb{Q}x_{i},\quad F:=\bigoplus_{i=0}^{\infty}\wedge^{i}V.$
    There is a natural grading on
    $F=\bigoplus_{i=0}^{\infty}F_{i}$
    by associating every $x_{i}$ to degree $1$. We consider the degree $1$ operator $K_{1}$ maps
    $x_{j_{1}}\wedge\cdots\wedge x_{j_{k}}\to x_{0}\wedge x_{j_{1}+1}\wedge\cdots\wedge x_{j_{k}+1};$
    Then $K_{1}$ is injective and induces a stable limit $F_{\infty}:=\lim_{i\to\infty}F_{i}$. A careful study of $F_{\infty}$ will reveal that
    $F_{\infty}\cong B.$
  2. (2)
    Let $p:\hat{S}\to S$ be the projection. Nakajima-Yoshioka [22] considered the moduli space of perverse coherent sheaves on $\hat{S}$, which connects $S^{[n]}$ and $\hat{S}^{[n]}$ by the following properties:
    1. (a)
      fixing a chern character in $H^{*}(\hat{S},\mathbb{Q})$
      $c_{n}=[\hat{S}]+n[pts]$
      there exists moduli space $M^{m}(c_{n})$ of perverse coherent sheaves for any integer $m\geq 0$ such that
      $M^{0}(c_{n})\cong S^{[n]},\quad M^{r}(c_{n})\cong\hat{S}^{[n]}\text{ if }r\gg 0.$
    2. (b)
      $M^{r}(c_{n})$ and $M^{r+1}(c_{n})$ are birational and connected through a wall-crossing.
    Thus $\coprod_{n=0}^{\infty}\hat{S}^{[n]}$ can be regarded as a stable limit of
    $\coprod_{n=0}^{\infty}M^{m}(c_{n}).$
Theorem 1.1 follows from the geometric representation theory of perverse coherent sheaves:
Theorem 1.3 (Theorem 6.2).
We have an identification
(1.4) $\bigoplus_{m=0}^{\infty}\bigoplus_{n=0}^{\infty}\mathbb{H}^{*}(M^{m}(c_{n}),\mathbb{Q})\cong\bigoplus_{n=0}^{\infty}\mathbb{H}^{*}(S^{[n]},\mathbb{Q})\otimes_{\mathbb{Q}}F$
and the action of the infinite dimensional Clifford algebra on $F$ is given by geometric correspondences.
1.3. A categorical Boson-Fermion correspondence
The fact that the bosonic Fock space $B$ is the stable limit of the Fermionic Fock space $F$ is a special case of a categorical Boson-Fermion correspondence which is of independent interest. It reveals a equivalence of certain infinite dimensional Heisenberg algebra (which we denote as $\mathcal{H}$) representations and graded infinte dimensional Clifford algebra (which we denote as $\mathcal{C}$) representations. The bridge between $\mathcal{C}$ and $\mathcal{H}$ is an $\mathbb{Q}$-algebra, which we call $\mathcal{E}$:
Definition 1.4 (2.2).
The algebra $\mathcal{E}$ is defined as a $\mathbb{Q}$-algebra generated by $K_{1},K_{-1},E,F$ with relations
$K_{-1}K_{1}=FE=1,\quad K_{-1}E=FK_{1}=0,\quad K_{1}K_{-1}+EF=1.$
The relation between $\mathcal{E}$ with $\mathcal{C}$ is direct: we consider $\{P_{i},Q_{i}\}_{i\in\mathbb{Z}_{>0}}$ such that $P_{1}=EK_{-1},Q_{1}=K_{1}F$ and
$P_{i}=K_{1}P_{i-1}K_{-1}-EP_{i-1}F,\quad Q_{i}=K_{1}Q_{i-1}K_{-1}-EQ_{i-1}F,\quad\text{ if }i>1.$
Then $P_{i},Q_{i}$ satisfy the relations of infinite dimensional Clifford algebras (see Proposition 2.2 for the details). On the other hand, the Fermionic Fock space $F$ can also be regarded as the highest weight representation of $\mathcal{E}$ generated by a vector $|0\rangle$ and the relations (see Proposition 2.8 for the details)
$F|0\rangle=0,\quad K_{-1}|0\rangle=K_{1}|0\rangle=|0\rangle.$
The relation between the representation theory of $\mathcal{E}$ and $\mathcal{H}$ is more delicate: roughly speaking, given a graded $\mathcal{E}$-representation on
$W=\bigoplus_{i\in\mathbb{Z}}W_{i}$
such that $K_{1}$ has degree $1$, then $K_{1}$ is injective and induces a stable limit
$W_{\infty}=\lim_{n\to\infty}W_{n}$
which is endowed with an $\mathcal{H}$ action if certain admissible property is satisfied. Particularly, we have the following equivalence
Theorem 1.5 (Categorical Boson-Fermion correspondence, Theorem 2.21).
We have a canonical equivalence between the following categories:
  1. (1)
    the abelian category of admissible $\mathcal{H}$-modules;
  2. (2)
    the abelian category of graded admissible $\mathcal{C}$-modules;
  3. (3)
    the abelian category of graded admissible $\mathcal{E}$-modules;
  4. (4)
    the abelian category of $\mathbb{Q}$-vector spaces.
Here the definitions of admissible modules are defined in Definitions 2.9, 2.10 and 2.13.
1.4. Derived Grassmannians and incidence varieties
By the categorical Boson-Fermion correspondence, what we would like to construct is an $\mathcal{E}$-module structure on
$\bigoplus_{m=0}^{\infty}\bigoplus_{n=0}^{\infty}\mathbb{H}^{*}(M^{m}(c_{n}),\mathbb{Q}).$
by incidence varieties. It follows from the incidence variety of the derived Grassmannian which is introduced by Jiang [11]: given a locally free sheaf $\mathcal{I}$ over a variety $X$, the Grassmannian (under the notation of Grothendieck) $\mathcal{I}$, which we denotes as $Gr(\mathcal{I},r)$, is the moduli space of quotients
$\mathcal{I}\twoheadrightarrow\mathcal{E}$
where $\mathcal{E}$ is a rank $r$ locally free sheaf on $X$. Jiang [11] generalized the concept of the Grassmannian and allows that $\mathcal{I}$ to be a perfect complex, i.e. locally a complex of locally free sheaves. The derived Grassmannian has good properties when $\mathcal{I}$ has tor-amplitude $[0,1]$, i.e. it is locally a two-term complex of locally free sheaves. Particularly, in this situation $\mathcal{I}^{\vee}[1]$ is also a tor-amplitude complex. For example, if $\mathcal{I}$ is the two term complex $V\to W$, then $\mathcal{I}^{\vee}[1]$ is $W^{\vee}\to V^{\vee}$.
Fixing integers $d_{-},d_{+}\geq 0$, we consider the derived Grassmannians
$Gr(\mathcal{I},d_{+}),\quad Gr(\mathcal{I}^{\vee}[1],d_{-}).$
In [11], Jiang defined an incidence variety
$Incid_{X}(\mathcal{I},d_{+},d_{-})$
which is the fiber product of $Gr(\mathcal{I},d_{+})$ and $Gr(\mathcal{I}^{\vee}[1],d_{-})$ over $X$ schematically but with a different perfect obstruction theory structure (i.e. derived structure in derived algebraic geometry). The (virtual) fundamental class of $Incid_{X}(\mathcal{I},d_{+},d_{-})$ induces correspondences:
$\phi_{d_{+},d_{-}}:\mathbb{H}^{*}(Gr(\mathcal{I}^{\vee}[1],d_{-}),\mathbb{Q})\to\mathbb{H}^{*}(Gr(\mathcal{I},d_{+}),\mathbb{Q}),$
$\psi_{d_{+},d_{-}}:\mathbb{H}^{*}(Gr(\mathcal{I},d_{+}),\mathbb{Q})\to\mathbb{H}^{*}(Gr(\mathcal{I}^{\vee}[1],d_{-}),\mathbb{Q}).$
Now we come back to the moduli space of perverse coherent sheaves $M^{m}(c_{n})$. Let $\mathcal{I}$ be the universal sheaf over $S^{[n]}\times S$ and
$\mathcal{I}_{o}:=\mathcal{I}|_{S^{[n]}\times o}.$
Then by Nakajima-Yoshioka [22], we have
$Gr_{S^{[n+\frac{l+l^{2}}{2}]}}(\mathcal{I}_{o}^{\vee}[1],l)\cong M^{l}(c_{n})\cong Gr_{S^{[n+\frac{-l+l^{2}}{2}]}}(\mathcal{I}_{o},l).$
Hence the indicence variety of derived Grassmannians induce the maps
$\psi_{l,l-1}:\bigoplus_{n\geq 0}\mathbb{H}^{*}(M^{l}(c_{n}),\mathbb{Q})\to\bigoplus_{n\geq 0}\mathbb{H}^{*}(M^{l-1}(c_{n}),\mathbb{Q}),$
$\phi_{l-1,l}:\bigoplus_{n\geq 0}\mathbb{H}^{*}(M^{l-1}(c_{n}),\mathbb{Q})\to\bigoplus_{n\geq 0}\mathbb{H}^{*}(M^{l}(c_{n}),\mathbb{Q}),$
$\psi_{l,l}:\bigoplus_{n\geq 0}\mathbb{H}^{*}(M^{l}(c_{n}),\mathbb{Q})\to\bigoplus_{n\geq 0}\mathbb{H}^{*}(M^{l}(c_{n-l}),\mathbb{Q}),$
$\phi_{l,l}:\bigoplus_{n\geq 0}\mathbb{H}^{*}(M^{l}(c_{n-l}),\mathbb{Q})\to\bigoplus_{n\geq 0}\mathbb{H}^{*}(M^{l}(c_{n}),\mathbb{Q}).$
The $\mathcal{E}$-representation on the cohomology of moduli space of perverse coherent sheaves follows from the above operators:
Theorem 1.6 (Theorem 6.1).
The operators
$E:=\bigoplus_{l>0}\phi_{l-1,l},\quad F:=\bigoplus_{l>0}\psi_{l-1,l},\quad K_{1}:=\bigoplus_{l\geq 0}\phi_{l,l},\quad K_{-1}:=\bigoplus_{l\geq 0}\psi_{l,l}$
forms an graded admissible $\mathcal{E}$-representation on
$\bigoplus_{m=0}^{\infty}\bigoplus_{n=0}^{\infty}\mathbb{H}^{*}(M^{m}(c_{n}),\mathbb{Q}).$
1.5. Some related work
1.5.1. The affine super-Yangian of $gl(1|1)$
Although not explicitly pointing it our paper, the representation we constructed is closed related to the quiver Yangian action of Li-Yamazaki [18] and Galakhov-Li-Yamazaki [8]. Here we take an example that $S$ is the equivariant $\mathbb{C}^{2}$ (although it is not projective). Li-Yamazaki [18] defined the affine super-Yangian of $gl(1|1)$, which we denote as
$Y_{\hbar_{1},\hbar_{2}}(\widehat{gl(1|1)}),$
as the $\mathbb{Q}(\hbar_{1},\hbar_{2})$ algebra generated by $e_{n}^{\circ/\bullet}$, $f_{n}^{\circ/\bullet}$, $\psi_{k}^{\circ/\bullet}$, $n\in\mathbb{Z}_{\geq 0}$, $k\in\mathbb{Z}$ with relations:
(1.5) $\begin{array}[]{l}\left\{e^{\circ}_{n},e^{\circ}_{k}\right\}=\left\{f^{\circ}_{n},f^{\circ}_{k}\right\}=\left[\psi^{\circ}_{n},e^{\circ}_{k}\right]=\left[\psi^{\circ}_{n},f^{\circ}_{k}\right]=\left\{e^{\bullet}_{n},e^{\bullet}_{k}\right\}=\left\{f^{\bullet}_{n},f^{\bullet}_{k}\right\}=\left[\psi^{\bullet}_{n},e^{\bullet}_{k}\right]=\left[\psi^{\bullet}_{n},f^{\bullet}_{k}\right]=0\,,\\ \left\{e_{n+2}^{\circ},e_{k}^{\bullet}\right\}-2\left\{e_{n+1}^{\circ},e_{k+1}^{\bullet}\right\}+\left\{e_{n}^{\circ},e_{k+2}^{\bullet}\right\}-\frac{\hbar_{1}^{2}+\hbar_{2}^{2}}{2}\left\{e_{n}^{\circ},e_{k}^{\bullet}\right\}+\frac{\hbar_{1}^{2}-\hbar_{2}^{2}}{2}\left[e_{n}^{\circ},e_{k}^{\bullet}\right]=0\,,\\ \left\{f_{n+2}^{\circ},f_{k}^{\bullet}\right\}-2\left\{f_{n+1}^{\circ},f_{k+1}^{\bullet}\right\}+\left\{f_{n}^{\circ},f_{k+2}^{\bullet}\right\}-\frac{\hbar_{1}^{2}+\hbar_{2}^{2}}{2}\left\{f_{n}^{\circ},f_{k}^{\bullet}\right\}-\frac{\hbar_{1}^{2}-\hbar_{2}^{2}}{2}\left[f_{n}^{\circ},f_{k}^{\bullet}\right]=0\,,\\ \left[\psi_{n+2}^{\circ},e_{k}^{\bullet}\right]-2\left[\psi_{n+1}^{\circ},e_{k+1}^{\bullet}\right]+\left[\psi_{n}^{\circ},e_{k+2}^{\bullet}\right]-\frac{\hbar_{1}^{2}+\hbar_{2}^{2}}{2}\left[\psi_{n}^{\circ},e_{k}^{\bullet}\right]+\frac{\hbar_{1}^{2}-\hbar_{2}^{2}}{2}\left\{\psi_{n}^{\circ},e_{k}^{\bullet}\right\}=0\,,\\ \left[\psi_{n+2}^{\circ},f_{k}^{\bullet}\right]-2\left[\psi_{n+1}^{\circ},f_{k+1}^{\bullet}\right]+\left[\psi_{n}^{\circ},f_{k+2}^{\bullet}\right]-\frac{\hbar_{1}^{2}+\hbar_{2}^{2}}{2}\left[\psi_{n}^{\circ},f_{k}^{\bullet}\right]-\frac{\hbar_{1}^{2}-\hbar_{2}^{2}}{2}\left\{\psi_{n}^{\circ},f_{k}^{\bullet}\right\}=0\,,\\ \left[\psi_{n+2}^{\bullet},e^{\circ}_{k}\right]-2\left[\psi_{n+1}^{\bullet},e_{k+1}^{\circ}\right]+\left[\psi_{n}^{\bullet},e_{k+2}^{\circ}\right]-\frac{\hbar_{1}^{2}+\hbar_{2}^{2}}{2}\left[\psi_{n}^{\bullet},e_{k}^{\circ}\right]-\frac{\hbar_{1}^{2}-\hbar_{2}^{2}}{2}\left\{\psi_{n}^{\bullet},e_{k}^{\circ}\right\}=0\,,\\ \left[\psi_{n+2}^{\bullet},f_{k}^{\circ}\right]-2\left[\psi_{n+1}^{\bullet},f_{k+1}^{\circ}\right]+\left[\psi_{n}^{\bullet},f_{k+2}^{\circ}\right]-\frac{\hbar_{1}^{2}+\hbar_{2}^{2}}{2}\left[\psi_{n}^{\bullet},f_{k}^{\circ}\right]+\frac{\hbar_{1}^{2}-\hbar_{2}^{2}}{2}\left\{\psi_{n}^{\bullet},f_{k}^{\circ}\right\}=0\,,\\ \left\{e^{\circ}_{n},f^{\circ}_{k}\right\}=-\psi^{\circ}_{n+k},\quad\left\{e^{\bullet}_{n},f^{\bullet}_{k}\right\}=-\psi^{\bullet}_{n+k},\quad\left\{e^{\circ}_{n},f^{\bullet}_{k}\right\}=\left\{e^{\bullet}_{n},f^{\circ}_{k}\right\}=0\,,\\ \left[\psi_{n}^{\circ/\bullet},\psi_{k}^{\circ/\bullet}\right]=0\,.\end{array}$
and higher order relations which we will discuss in our future work.
When $S$ is the equivariant $\mathbb{C}^{2}$, the moduli space of perverse coherent sheaves was described as a quiver-like variety by Nakajima-Yoshioka [23]. The affine super-Yangian $Y_{\hbar_{1},\hbar{2}}(\widehat{gl(1|1)})$ acts on the cohomology of moduli spaces of perverse coherent sheaves by Galakhov-Morozov-Tselousov [9]. The operators $\phi_{l-1,l},\phi_{l,l},\psi_{l-1,l},\psi_{l,l}$ in Theorem 1.6 actually comes from the action of certain elements in $Y_{\hbar_{1},\hbar_{2}}(\widehat{gl(1|1)})$, which we will discuss in the future work.
1.5.2. Cohomological Hall algebra (COHA)
An equivalent definition of the affine super-Yangian $Y_{\hbar_{1},\hbar_{2}}(\widehat{gl(1|1)})$ is the double COHA on the resolved conifold by Soibelmann-Rapcak-Yang-Zhao [26]. Actually, our paper should be regarded as a continuation of Section 7.2.2 of their paper.
By the dimension reduction, the affine super-Yangian is also the double COHA of certain one-dimensional objects over a surface, which was defined by Diaconescu-Porta-Sala [6]. A systematically computation of the double COHA will appear very soon [7].
1.5.3. Koseki’s categorification
Another work which we get great inspiration is the categorification of the derived category of coherent sheaves on the Hilbert scheme of points [15] by Koseki. To get more precise representation theoretic relations, we apply the semi-orthogonal decomposition theorem of Jiang [13] with explicit Fourier-Mukai kernels.
1.6. Notations in this paper
[…]
All the varieties and derived schemes in this paper will be over $\mathbb{C}$.
1.7. Acknowledgements
I would like to thank Hiraku Nakajima, Andrei Neguţ, Francesco Sala, Qingyuan Jiang, Jiakang Bao, Dmitry Galakhov and Henry Liu for many interesting discussions on the subject.
2. A categorical Boson-Fermion correspondence
Consider the infinite-dimensional Clifford algebra $\mathcal{C}$ generated by $\{P_{i},Q_{i}\}_{i\in\mathbb{Z}_{>0}}$ satisfying the anti-commutator relations
(2.1) $\{P_{i},P_{j}\}=0,\quad\{Q_{i},Q_{j}\}=0,\quad\{P_{i},Q_{j}\}=\delta_{ij}$
and the infinite-dimensional Heisenberg algebra $\mathcal{H}$ generated by $\{H_{i}\}_{i\in\mathbb{Z}-\{0\}}$ with the commutator relations
$[H_{i},H_{j}]=i\delta_{i,-j}$
In this section, we demonstrate an equivalence between specific representations of $\mathcal{C}$ and $\mathcal{H}$ which can be regarded as a categorical version of a Boson-Fermion correspondence.
The equivalence is established through a new algebra, denoted as $\mathcal{E}$, generated by $K_{1},K_{-1},E,F$ with the relations
(2.2) $K_{-1}K_{1}=FE=1,\quad K_{-1}E=FK_{1}=0,\quad K_{1}K_{-1}+EF=1.$
The algebra $\mathcal{E}$ naturally emerges from the (semi)-orthogonal decomposition of vector spaces or linear categories: given a decomposition of a vector space
$V\cong V_{0}\oplus V_{1},$
let $F$ (resp. $E$) be the projection (resp. embedding) from $V$ to $V_{1}$ (resp. from $V_{1}$ to $V$). Let $K_{-1}$ (resp. $K_{1}$) be the projection (resp. embedding) from $V$ to $V_{0}$ (resp. $V_{0}$ to $V$). The operators satisfy the relations 2.2.
2.1. Basic properties of $\mathcal{E}$ and $\mathcal{E}$-representations
Let
$E_{i}=K_{1}^{i}E,\quad F_{i}=FK_{-1}^{i},\quad i\geq 0.$
Through straightforward calculations, we find that
Lemma 2.1.
The operators $E_{i},F_{i}$ satisfy
(2.3) $\sum_{i=0}^{l}E_{i}F_{i}=1-K_{1}^{l+1}K_{-1}^{l+1},\quad E_{i}F_{i}=K_{1}^{i}K_{-1}^{i}-K_{1}^{i+1}K_{-1}^{i+1};$
and
$K_{1}E_{i}=E_{i+1},\quad K_{-1}E_{i+1}=E_{i},\quad$ $F_{i}K_{-1}=F_{i+1},\quad F_{i+1}K_{1}=F_{i},$
$F_{i}E_{j}=\delta_{i,j},\quad i,j\geq 0$
Let $P_{1}=EK_{-1},Q_{1}=K_{1}F$ and
$P_{i}=K_{1}P_{i-1}K_{-1}-EP_{i-1}F,\quad Q_{i}=K_{1}Q_{i-1}K_{-1}-EQ_{i-1}F,\quad\text{ if }i>1.$
in Section A.1 we prove that
Proposition 2.2.
The elements $\{P_{i},Q_{i}\}_{i\in\mathbb{Z}_{\geq 0}}$ satisfy the relations of $\mathcal{C}$, i.e. 2.1.
Lemma 2.3.
Given a $\mathcal{E}$-module $W$, the action of $K_{1},E$ on $W$ are injective and the action of $K_{-1},F$ on $W$ are surjective. Moreover,
$K_{1}W\cong ker(F),\quad EW\cong ker(K_{-1}),\quad W\cong K_{1}W\oplus EW.$
as vector spaces.
Proof.
The injective (resp. surjective) property of $K_{1},E$ (resp. $K_{-1},F$) follows from
$K_{-1}K_{1}=FE=1.$
Now we prove that $K_{1}W\cong ker(F)$ (and similarly $EW\cong Ker(K_{-1})$). First we have $K_{1}W\subset ker(F)$ as $FK_{1}=0$. Second, for any $f\in Ker(F)$, we have
$f=K_{1}K_{-1}f+EFf=K_{1}(K_{-1}f)\in K_{1}W.$
Moreover, if $K_{-1}f=0$, then $f=0$ and hence
$ker(f_{0})\cap ker(f_{1})=K_{1}W\cap EW=0$
and thus we have $W\cong K_{1}W\bigoplus EW$. ∎
Remark 2.4.
Any non-zero $\mathcal{E}$-module has to be infinite-dimensional: for a finite-dimensional representation, we have
$K_{1}K_{-1}=id,\quad EF=id,$
which contradicts with that $K_{1}K_{-1}+EF=id$.
2.2. Highest representation theory of $\mathcal{E}$ and Fermionic Fock space
Definition 2.5.
Given $\lambda\in\mathbb{Q}^{*}$, an $\mathcal{E}$-module $W$ is defined to be the highest weight representation of weight $\lambda$ if there exists $w\in W$ which generates $W$ and
$Fw=0,\quad K_{1}w=\lambda w.$
The verma module of the highest weight representation of weight $\lambda$ is denoted as $F_{\lambda,\mathcal{E}}$, and the hightest weight vector is denoted as 00$\rangle$0 ⟩. We denote $F_{\mathcal{E}}:=F_{1,\mathcal{E}}$.
Given $\lambda\in\mathbb{Q}^{*}$, there is an automorphism of $\mathcal{E}$ which maps
$E\to\lambda E,\quad K_{1}\to\lambda K_{1},\quad F\to\lambda F,\quad K_{-1}\to\lambda K_{-1},$
and under this auto equivalence, the representation $F_{\lambda,\mathcal{E}}$ is mapping to $F_{\mathcal{E}}$.
Now we explain that $F_{\mathcal{E}}$ is a Fermionic Fock space: given $i\in\mathbb{Z}_{>0}$, we denote $x_{i}$ as a formal symbol. We consider the infinite-dimensional vector space
$V=\bigoplus_{i=1}^{\infty}\mathbb{Q}x_{i}.$
Definition 2.6.
The Fermionic Fock space $F$ is defined as the infinite wedge product
$F:=\bigoplus_{i=1}^{\infty}\wedge^{i}F$
and it has basis
$x_{i_{1}}\wedge\cdots\wedge x_{i_{k}}$
for any sequence $0<i_{1}<i_{2}<\cdots<i_{k},k\geq 0$.
Through straightforward calculations, we find that
Lemma 2.7.
There is a representation of $\mathcal{E}$ on the Fermionic Fock space $F$, such that the action of $id$ is identity and
$K_{1}(x_{j_{1}}\wedge\cdots\wedge x_{j_{k}})=x_{j_{1}+1}\wedge\cdots\wedge x_{j_{k}+1},$
$K_{-1}(x_{j_{1}}\wedge\cdots\wedge x_{j_{k}})=x_{j_{1}-1}\wedge\cdots\wedge x_{j_{k}-1},$
$E(x_{j_{1}}\wedge\cdots\wedge x_{j_{k}})=x_{1}\wedge x_{j_{1}+1}\wedge\cdots\wedge x_{j_{k}+1},$
$F(x_{j_{1}}\wedge\cdots\wedge x_{j_{k}})=\frac{\partial}{\partial x_{1}}(x_{j_{1}-1}\wedge\cdots\wedge x_{j_{k}-1}).$
where $x_{0}:=0$. Under the action of $\mathcal{E}$ on $F$, the action of $P_{i}$ is wedging $x_{i}$ and the action of $Q_{i}$ is $\frac{\partial}{\partial x_{i}}$.
Proposition 2.8.
We have
$F_{\mathcal{E}}\cong F$
which is a irreducible representation.
Proof.
Let $|0\rangle$ be the highest weight vector of $F_{\mathcal{E}}$. Then themap
$|0\rangle\to 1$
induces a morphism from $F_{\mathcal{E}}$ to $F$. We notice that all the vectors
$E_{i_{1}}E_{i_{2}-i_{1}}\cdots E_{i_{n}-i_{n-1}}|0\rangle$
where $0\leq i_{1}\leq i_{2}\leq\cdots\leq i_{n}$ forms a basis of $F_{\mathcal{E}}$, and its image is
$x_{i_{1}}\wedge x_{i_{2}+1}\wedge\cdots\wedge x_{i_{n}+n}.$
Hence the map from $F_{\mathcal{E}}$ to $F$ is an isomorphism.
Now we prove that the Fermionic Fock space $F$ is irreducible as a $\mathcal{E}$-module. Let $a=\oplus_{i=1}^{N}a_{i}\in F$, we take a monomial
$x_{j_{1}}\wedge x_{j_{2}}\wedge\cdots x_{j_{N}}$
in $a_{N}$ and assume its coefficient is $1$. Then
$F_{j_{N}-j_{N-1}-1}\cdots F_{j_{1}-1}a=1.$
Thus the module $\mathcal{E}$ is irreducible. ∎
2.3. Admissible representations and Boson-Fermion correspondence
We consider the category of graded $\mathcal{E}$-modules (resp. $\mathcal{C}$-repsentations), i.e. $\mathcal{E}$-modules (resp. $\mathcal{C}$-modules)
$W\cong\bigoplus_{i\in\mathbb{Z}}W_{i}$
such that the degree of $E,F,K_{\pm 1}$ are $1,-1,0$ respectively (resp. the degree of $P_{i}$ is $1$ and $Q_{i}$ is $-1$).
Definition 2.9.
A graded $\mathcal{E}$-module
$W\cong\bigoplus_{i=0}^{\infty}W_{i}$
is called admissible if
  1. (1)
    $K_{1}|_{W_{0}}=K_{-1}|_{W_{0}}=id$;
  2. (2)
    for any $x\in W_{i}$ such that $i>0$, there exists $N>0$ such that $K_{-1}^{N}x=0$.
Definition 2.10.
A graded $\mathcal{C}$-module
$W:=\bigoplus_{i\geq 0}W_{i}$
is called admissible if
  1. (1)
    for any $x\in W$, there exists $N>0$ such that $Q_{m}x=0$ when $m>0$.
  2. (2)
    We have
    $W_{0}\cong\bigcap_{i>0}ker(Q_{i}).$
Example 2.11.
On the Fermionic Fock space, we assoiciate $x_{i}$ with degree $1$. Then $F$ is a graded admissible $\mathcal{E}$-module.
Example 2.12.
We consider the Bosonic Fock space
$B:=\bigoplus_{i=0}^{\infty}V\cong\mathbb{Q}[y_{1},y_{2},\cdots].$
It has the $\mathcal{H}$-representation structure by
$\begin{cases}H_{k}\to y_{k},\quad k>0,\\ H_{k}\to\frac{\partial}{\partial y_{-k}},\quad k<0\end{cases}$
which is an admissible representation.
We also introduce the admissible $\mathcal{H}$-representations:
Definition 2.13.
Given a sequence of non-negative integers $\lambda=(n_{1},n_{2},\cdots)$ such that $n_{m}=0$ when $m\gg 0$, we denote
$|\lambda|=\sum_{i=1}^{\infty}in_{i},\quad H_{\lambda}=\prod_{i=1}^{\infty}H_{i}^{n_{i}},\quad H_{-\lambda}=\prod_{i=1}^{\infty}H_{-i}^{n_{i}}.$
An $\mathcal{H}$-representation $W$ is called admissible, if for any $x\in W$ and $i>0$, there exists an integer $N>0$ such that $H_{-\lambda}x=0$ for any $|\lambda|>N$.
In Sections 2.4, 2.5 and 2.6, we establish the following equivalence
Proposition 2.14.
We have an equivalence among the abelian category of admissible $\mathcal{H}$-modules, graded admissible $\mathcal{C}$-modules and graded admissible $\mathcal{E}$-modules.
2.4. Graded $\mathcal{E}$-modules
We first establish the equivalence between the category admissible graded $\mathcal{E}$-modules and graded $\mathcal{C}$-modules. For any $\mathcal{E}$-module $W$, the morphsim from $\mathcal{C}$ to $\mathcal{E}$ induces a $\mathcal{C}$-representation on $W$.
Proposition 2.15.
If $W$ is an admissible graded $\mathcal{E}$-module, the induced $\mathcal{C}$-module on $W$ is also admissible.
Proof.
We only need to prove that
$W_{0}\cong\bigcap_{i>0}ker(Q_{i}).$
First we notice that for any non-zero $x\in\oplus_{i>0}W_{i}$, there exists $M\geq 0$ such $F_{M}x\neq 0$. Otherwise, by 2.3, for any $M>0$
$K_{-1}^{M}x=x$
which contradicts with the assumption of admissibility. Let $M_{0}$ be the smallest integer such that $F_{M_{0}}x\neq 0$. Then by induction, we have
$Q_{M+1}x=K_{1}Q_{M}K_{-1}x=\cdots=K_{1}^{M_{0}}Q_{1}K_{-1}^{M_{0}}x=K_{1}^{M_{0}+1}F_{M}x.$
which is non-zero as $K_{1}$ is injective. ∎
Let $W$ be an admissible $\mathcal{C}$-module. Let
$P(z):=\sum_{i=1}^{\infty}P_{i}z^{i},\quad Q(z):=\sum_{i=1}^{\infty}Q_{i}z^{i}$
For any operator $T$ on $W$, we consider
$\Gamma(T)(z):=\sum_{i\in\mathbb{Z}}\Gamma_{i}(T)z^{i}$
such that
$\Gamma_{i}(T)=\sum_{\begin{subarray}{c}m-n=i\\ m,n\geq 1\end{subarray}}P_{m}TQ_{n}$
which is well defined due to the admissible property. We also formally written
$\Gamma(T)(z):=P(z)TQ(z^{-1}).$
We denote
$\Lambda:=\Gamma_{0}(id).$
by Lemma A.4 we have $\Lambda|_{W_{i}}=i\cdot id$ and hence is a invertible operator.
By induction there are unique operators $K_{1}$, $K_{-1}$ on $W$ satisfy the following relations
  1. (1)
    $K_{1},K_{-1}$ are identity on $W_{0}$;
  2. (2)
    we have
    $\Lambda K_{1}=\Gamma_{1}(K_{1}),\quad\Lambda K_{-1}=\Gamma_{-1}(K_{-1}).$
We also define
$E:=P_{1}K_{1},\quad F:=K_{-1}Q_{1}.$
In Section A.2, we will prove that
Proposition 2.16.
The operators $K_{1},K_{-1},E,F$ induce an admissible $\mathcal{E}$-module on $W$.
2.5. Filtration and Fermionization
In this subsection we give the functor from admissible $\mathcal{H}$-representations to graded admissible $\mathcal{E}$-representations: let $W_{\infty}$ be an admissible $\mathcal{H}$-representation, we consider
$W_{i}:=\bigcap_{l>i}ker(H_{l}).$
It satisies that $W_{i}\subset W_{j}$ for $i\leq j$ and
$W_{\infty}=\bigcup_{i\geq 0}W_{i}.$
Hence it forms a filtration of $W$. For any $k>0$, we consider operators
$\mathcal{L}_{k}:=\sum_{i=1}^{\infty}\frac{-1}{k^{i}i!}H_{k}^{i-1}H_{-k}^{i}$
The admissible property of $W_{\infty}$ makes sure that the action of $\mathcal{L}_{k}$ on $W_{\infty}$ is well defined. By the commutative property of the Heisenberg algebra, we have
(2.4) $H_{k}W_{k}\subset W_{k},\quad\mathcal{L}_{k}W_{k}\subset W_{k}.$
In Section A.3, we will prove that
Proposition 2.17.
We have
(2.5) $(Id-H_{k}\mathcal{L}_{k})W_{k}\subset W_{k-1}$
and thus the following operators
$E=\bigoplus_{i\geq 0}id,\quad F=\bigoplus_{i\geq 0}(id-H_{k}\mathcal{L}_{k})$
are well defined on
$W=\bigoplus_{i\geq 0}W_{i}.$
Moreover, considering the operators
$K_{1}:=\bigoplus_{k\geq 0}H_{k},\quad K_{-1}:=\bigoplus_{k\geq 0}\mathcal{L}_{k},$
where $H_{0}=\mathcal{L}_{0}:=id$. Then $W$ is an graded admissible $\mathcal{E}$-module.
2.6. Stable limit and Bosonization
The functor from graded admissible $\mathcal{E}$-modules to admissible $\mathcal{H}$-modules is constructed by taking the stable limit:
Definition 2.18.
Given a graded admissible $\mathcal{E}$-module
$W:=\bigoplus_{i\geq 0}W_{i}$
we define the stable limit, denotes $W_{\infty}$ as the limit
$\lim_{n\to\infty}W_{i}$
where the morphism form $W_{i}$ to $W_{i+1}$ is $K_{1}$ which is injective.
By induction, there exists a unique series of operators $\{H_{i,j}\}|_{-i\leq j\leq i,j\neq 0}$ which acts on $W_{i}$ such that
$H_{i,i}:=K_{1}|_{W_{i}},\quad H_{-i,i}:=-i\sum_{l=0}^{\infty}(K_{1})^{l}K_{-1}^{l+1}|_{W_{i}}.$
and if $|j|\neq i$
$H_{i,j}:=\sum_{l=0}^{\infty}E_{l}H_{i-1,j}F_{l}.$
Moreover, as
$H_{i+1,j}E=\sum_{l=0}^{\infty}E_{l}H_{i,j}F_{l}E=\sum_{l=0}^{\infty}E_{l}H_{i,j}\delta_{0,l}=E_{l}H_{i,j},$
for any $|j|\leq i$, the operators $\{H_{i,l}\}_{i\geq 0}$ converges to an operator $H_{l}$ on $W_{\infty}$. In the Section A.4, we will prove that
Proposition 2.19.
The operators $H_{l}|_{l\in\mathbb{Z}-\{0\}}$ forms an admissible $\mathcal{H}$-representation on $W_{\infty}$.
Proof of Proposition 2.14.
It follows from Propositions 2.15, 2.16, 2.17 and 2.19. ∎
2.7. The Boson-Fermion correspondence
The Fermion Fock space $F$ is a graded admissible $\mathcal{E}$-module and the Bosonic Fock space $B$ is an admissible $\mathbb{H}$-module. We have that
Proposition 2.20.
The stable limit of the Fermionic Fock space $F$ is the Bosonic Fock space $B$.
Proof.
We construct the mapping from $\rho:F\to B$
$\rho(x_{i_{1}}\wedge x_{i_{2}}\cdots\wedge x_{i_{n}}):=y_{1}^{i_{n}-i_{n-1}-1}y_{2}^{i_{n-1}-i_{n-2}-1}\cdots y_{1}^{i_{1}-1}.$
It is easy to see the following properties of $\rho$
  1. (1)
    $\rho$ is invariant under $K_{1}$, i.e. for any $x\in F$,
    $\rho(K_{1}x)=\rho(x).$
  2. (2)
    For any $i\geq 0$, $\rho$ induces an isomorphism of vector spaces between $F_{i}$ and
    $B_{i}\cong\mathbb{Q}[y_{1},\cdots,y_{i}].$
hence by taking the limit $i\to\infty$, we prove Proposition 2.20. ∎
Now we consider the functor from the abelian category of vector spaces to the abelian category of admissible $\mathcal{H}$-representations
$\mathfrak{b}:W_{0}\to B\otimes_{\mathbb{Q}}W_{0}$
and the functor from the abelian category of vector spaces to the abelian category of graded admissible $\mathcal{E}$-representations
$\mathfrak{f}:W_{0}\to F\otimes_{\mathbb{Q}}W_{0}$
Theorem 2.21 (Categorical Boson-Fermion Correspondence).
The functors $\mathfrak{b}$ or $\mathfrak{f}$ is an equivalence between the category of vector spaces to the category of admissible $\mathcal{H}$-modules or the category of admissible $\mathcal{E}$ or $\mathcal{C}$-modules.
By Proposition 2.14, Theorem 2.21 follows from the following proposition:
Proposition 2.22.
For any admissible $\mathcal{H}$-module $W_{\infty}$, we have
$W_{\infty}\cong B\otimes_{\mathbb{Q}}W_{0}.$
Proof.
The morphism from $B\otimes_{\mathbb{Q}}W_{0}$ to $W_{\infty}$ by
$y_{1}^{l_{1}}\cdots y_{m}^{l_{m}}\otimes w_{0}\to H_{1}^{l_{1}}\cdots H_{m}^{l_{m}}w_{0}$
is a morphism of $\mathcal{H}$-modules. Now we prove that it is an isomorphism as vector spaces. By induction, we only need to prove that the morphism
$W_{i}\otimes_{\mathbb{Q}}\mathbb{Q}[y_{i+1}]$ $\to W_{i+1},$
$w_{0}+w_{1}y_{i+1}+\cdots+w_{n}y_{i+1}^{n}$ $\to w_{0}+H_{i+1}w_{1}+\cdots+H_{i+1}^{n}w_{1}.$
is an isomorphism. It is injective as given a monomial
$w=w_{0}+w_{1}y_{i+1}+\cdots+w_{n}y_{i+1}^{n}$
such that $w_{i}\in W_{i}$ and $w_{n}\neq 0$, we have $H_{-i-1}^{n}w=n!w_{n}$ and its image is non-zero.
To prove that it is surjective, let $w\in W_{i+1}$. We have
$w=\sum_{k=0}^{N}H_{i+1}^{k}((1-H_{i+1}\mathcal{L}_{i+1})\mathcal{L}_{i+1}^{k}w).$
for a sufficient large $N$ and for any $k\geq 0$, by 2.5
$(1-H_{i+1}\mathcal{L}_{i+1})\mathcal{L}_{i+1}^{k}w\in W_{i}.$
Hence the morphism is surjective. ∎
3. Derived categories of derived Grassmannians
Given a variety $X$, a locally free sheaf $\mathcal{V}$ and an integer $r\geq 0$, the Grassmannian (on Grothendieck’s notation) $Gr_{X}(\mathcal{V},r)$ is defined as the moduli space of quotients
$\mathcal{V}\twoheadrightarrow\mathcal{E}$
where $\mathcal{E}$ is a locally free sheaf of rank $r$.
Jiang [12, 13, 11] extended the definition of Grassmannians and flag varieties such that $\mathcal{V}$ can be any perfect complex. Particularly, when $\mathcal{V}$ is a tor-amplitude $[0,1]$ perfect complex, i.e. locally a two term complexes of locally free sheaves
$\mathcal{V}\cong\{W\to V\}$
the derived Grassmannian is quasi-smooth (i.e. the locally complete intersection scheme in the derived algebraic geometry setting). Moreover, Jiang [13] strengthen the result of Toda [28] for the semi-orthogonal decomposition of derived Grassmannians by given the Fourier-Mukai kernels.
The concept of derived Grassmannians and flag varieties included many important incidence varieties in geometric representation theory. We will refer to [25] and [20] for several examples. In this paper, we will focus on the example in Section 5, i.e. the moduli space of perverse coherent sheaves on the blow-ups introduced by Nakajima-Yoshioka [22].
Now we review several main results of Jiang [12, 13, 11]. We will adapt the language of derived algebraic geometry and $\infty$-categories in this section. The relation between derived geometry and (classical) birational geometry can be described through the degeneracy theory, which we explained in a companion paper [29].
In this section, we will always assume that $X$ is a derived scheme and $\mathcal{E}$ is a Tor-amplitude $[0,1]$ perfect complex on $X$.
3.1. Derived Grassmannians of a Tor-amplitude $[0,1]$ complex
Definition 3.1 (Definition 4.3 of [11]).
Given an integer $d\geq 0$, we define the derived Grassmannian
$Gr_{X}(\mathcal{E},d)$
as the derived stack which parametrized the moduli space of quotients
$\{\mathcal{E}\twoheadrightarrow\mathcal{V}\}$
where $\mathcal{V}$ is a rank $d$ locally free sheaf. We define
$pr_{\mathcal{E},d}:Gr_{X}(\mathcal{E},d)\to X$
as the projection morphism.
Remark 3.2.
Jiang’s definition requires that $d\geq 1$. Here we allow $d=0$ be defining
$Gr_{X}(\mathcal{E},0)\cong X.$
The concept of the quasi-smooth morphisms (schemes) is a generalization of the locally complete intersection morphisms (schemes) in derived algebraic geometry:
Definition 3.3.
A morphism of derived schemes $f:X\to Y$ is defined to be quasi-smooth if the cotangent complex $L_{f}:=L_{X/Y}$ is of tor-amplitude $[0,1]$. A derived scheme $X$ is defined to be quasi-smooth if the morphism from $X$ to $\mathrm{Spec}(\mathbb{C})$ is quasi-smooth. For a quasi-smooth morphism $f$ (resp. scheme $X$), we define its virtual dimension, denoted as
$vdim(f),\quad(\text{resp. }vdim(X))$
as the rank of the cotangent complex $L_{f}$ (resp. $L_{X}$). The relative canonical bundle $K_{f}$ (resp. $K_{X}$ )is defined as the determinant of $L_{f}$ (resp. $L_{X}$). (We refer to [27] or [12] for the determinant of a tor-amplitude $[0,1]$ complex).
Remark 3.4.
Given a quasi-smooth derived scheme $X$ over $\mathbb{C}$, we say that $X$ is classical if it has trivial derived structure, i.e. $\pi_{0}(X)\cong X$ where $\pi_{0}(X)$ is the underlying scheme. By Corollary 2.11 of [1] (which we also discussed in the Appendix A of [30]), $X$ is classical if and only if
$vdim(X)=dim(\pi_{0}(X)).$
and $X$ is always a locally complete intersection variety if $X$ is classical.
Theorem 3.5 ([11]).
The derived Grassmannian $Gr_{X}(\mathcal{E},d)$ is a proper derived scheme over $X$ and the projection map
$pr_{\mathcal{E},d}:Gr_{X}(\mathcal{E},d)\to X$
is quasi-smooth with relative virtual dimension is $d(rank(\mathcal{E})-d)$.
Let $v_{d}$ be the universal quotient map over $Gr_{X}(\mathcal{E},d)$
$pr_{\mathcal{E},d}^{*}\mathcal{E}\xrightarrow{v_{d}}\mathcal{V}_{d}$
Proposition 3.6 (Theorem 4.18 of [11]).
The cotangent complex
$L_{Gr_{X}(\mathcal{E},d)/X}\cong\mathcal{V}_{d}^{\vee}\otimes fib(v_{d})$
where $fib$ denotes the fiber (i.e. cylinder) of a map.
Corollary 3.7.
The relative canonical bundle of $pr_{\mathcal{E},d}$ is
$K_{Gr_{X}(\mathcal{E},d)}:=det(L_{Gr_{X}(\mathcal{E},d)/X})\cong pr_{\mathcal{E},d}^{*}det(\mathcal{E})\otimes det(\mathcal{V}_{d})^{-1}.$
For a proper quasi-smooth morphism $f:X\to Y$, the derived pullback and pushforward are all well-defined in the $\infty$-category of perfect complexes
$f^{*}:Perf(Y)\to Perf(X),\quad f_{*}:Perf(X)\to Perf(Y).$
Let $f_{!}$ be the left-adjoint functor $f^{*}$, and $f^{!}$ by the right-adjoint functor of $f_{*}$. By the Grothendieck duality, we have
(3.1) $f_{!}\cong f_{*}(-\otimes K_{Y/X})[vdim(f)],\quad f^{!}\cong(-\otimes K_{Y/X})\circ f^{*}[vdim(f)]$
where $vdim(f)$ is the rank of the cotangent complex $L_{f}$, which is a (locally) constant for a quasi-smooth morphism.
3.2. The incidence varieties
We notice that if $\mathcal{E}$ has tor-amplitude $[0,1]$, the shifted dual $\mathcal{E}^{\vee}[1]$ also has tor amplitude $[0,1]$.
Definition 3.8 (Incidence correspondence, Definition 2.7 of [13]).
Given $(d_{+},d_{-})\in\mathbb{Z}_{\geq 0}^{2}$, the incidence scheme $Inc_{X}(\mathcal{E},d_{+},d_{-})$ can be defined in two ways, which were proven to be equivalent in [11]:
  1. (1)
    Over the projection morphism $pr_{+}:Gr_{X}(\mathcal{E},d_{+})\to X$, let $\mathcal{E}_{+}$ be the fiber of the universal quotient
    $\theta_{+}:pr^{*}_{+}\mathcal{E}\to\mathcal{V}^{+}.$
    we define
    $Inc_{X}(\mathcal{E},d_{+},d_{-}):=Gr_{Gr_{X}(\mathcal{E},d+)}(\mathcal{E}_{+}^{\vee}[1],d_{-}).$
  2. (2)
    Over the projection morphism $pr_{-}:Gr_{X}(\mathcal{E}^{\vee}[1],d_{-})\to X$, let $\mathcal{E}_{-}$ be the fiber of the universal quotient
    $\theta_{-}:pr^{*}_{-}\mathcal{E}^{\vee}[1]\to\mathcal{V}^{-\vee}.$
    we define
    (3.2) $nc_{X}(\mathcal{E},d_{+},d_{-}):=Gr_{Gr_{X}(\mathcal{E}^{\vee}[1],d_{-})}(\mathcal{E}_{-}^{\vee}[1],d^{+}).$
The above definitions induce canonical projection morphisms:
$r^{+}_{d_{+},d_{-}}:Inc_{X}(\mathcal{E},d_{+},d_{-})\to Gr_{X}(\mathcal{E},d_{+}),\quad r^{-}_{d_{+},d_{-}}:Inc_{X}(\mathcal{E},d_{+},d_{-})\to Gr_{X}(\mathcal{E}^{\vee}[-1],d_{-}).$
and we will also make it short as $r^{+},r^{-}$ respectively if it does not make any confusion.
As $r^{+}$ and $r^{-}$ are both quasi-smooth and proper, the functor
$r^{+}_{*}r^{-*}:Perf(Gr_{X}(\mathcal{E}^{\vee}[-1],d_{-}))\to Perf(Gr_{X}(\mathcal{E},d_{+}))$
is represented by the Fourier-Mukai transform
$\Phi_{d_{+},d_{-}}:=(r^{+}\times r^{-})_{*}(\mathcal{O}_{Inc_{X,d_{+},d_{-}}})\in QCoh(Gr_{X}(\mathcal{E}^{\vee}[-1],d_{-})\times Gr_{X}(\mathcal{E},d_{+})),$
Here $QCoh$ denotes the $\infty$-category of quasi-coherent sheaves and $Perf$ denotes the $\infty$-category of perfect complexes.
Now we compute the left and right adjoint functor of $\Phi_{d_{+},d_{-}}$. Let $e$ the rank of $\mathcal{E}$. First we notice that the relative virtual dimension of $r^{+}$ and $r^{-}$ is
$vdim(r^{+})=d_{-}(-e+d_{+}-d_{-}),\quad vdim(r^{-})=d_{+}(e+d_{-}-d_{+}).$
and the virtual dimension of the incidence variety $Inc_{X}(\mathcal{E},d_{+},d_{-})$ over $X$ is
$d_{-}(-e-d_{-})+d_{+}(e-d_{+})+d_{-}d_{+}.$
Let $\Psi_{d_{+},d_{-}}^{L}$ and $\Psi_{d_{+},d_{-}}^{R}$ be the left and adjoint functor of $\Phi_{d_{+},d_{-}}$. By the Grothendieck duality 3.1, we have
$\Psi_{d_{+},d_{-}}^{L}\cong(r^{-}\times r^{+})_{*}(K_{r^{-}}[vdim(r^{-})]),\quad\Psi_{d_{+},d_{-}}^{R}\cong(r^{-}\times r^{+})_{*}(K_{r^{+}}[vdim(r^{+})])$
as the Fourier-Mukai kernels.
3.3. Semi-orthogonal decomposition and decomposition of the diagonal
When $e:=rank(\mathcal{E})\geq 0$ is positive, by Toda [28] and Jiang [13], for any $d\geq 0$ the category
$Perf(Gr_{X}(\mathcal{E},d))$
has a semi-orthogonal decomposition by copies of
$Perf(Gr_{X}(\mathcal{E}^{\vee}[1],d^{\prime})),\quad max\{d-e,0\}\leq d^{\prime}\leq d.$
Moreover, Jiang [13] proved that the functors are given as Fourier-Mukai transforms through the incidence varieties. Now we introduce a special case when the rank of $\mathcal{E}$ is $1$:
Theorem 3.9 ([13]).
If $e=1$, then for any $d>0$, the functors $\Phi_{d,d}$ and $\Phi_{d.d-1}$ are all fully faithful and we have the semi-orthogonal decomposition
$Perf(Gr_{X}(\mathcal{E},d))=<Im(\Phi_{d,d-1}),Im(\Phi_{d,d})>.$
Corollary 3.10.
Let $\Psi_{d,d-1}^{R}$ be the right adjoint functor of $\Phi_{d,d-1}$ and $\Psi_{d,d}^{L}$ be the left-adjoint fucntor of $\Phi_{d,d}$. Then there exsits a canonical triangle of Fourier-Mukai transforms
(3.3) $\Phi_{d,d-1}\Psi_{d,d-1}^{R}\to\mathcal{O}_{\Delta}\to\Phi_{d,d}\Psi_{d,d}^{L}$
where $\Delta$ is the diagonal embedding of $Gr_{X}(\mathcal{E},d)$. Moreover, we have
(3.4) $\Psi_{d,d-1}^{R}\Phi_{d,d-1}\cong\mathcal{O}_{\Delta},\quad\Psi_{d,d}^{L}\Phi_{d,d}\cong\mathcal{O}_{\Delta},\quad\Psi_{d,d-1}^{R}\Phi_{d,d}\cong 0,\quad\Psi_{d,d}^{L}\Phi_{d,d-1}\cong 0.$
Proof.
We only need to show natural transforms of functors are induced by morphisms of Fourier-Mukai transforms. As all the setting can be naturally lifted to $\infty$-categories, it follows from Theorem 1.2(2) of [3]. ∎
4. Mukai pairing and cohomological correspondences
In this section we always assume that $X$ and $Y$ be smooth projective varieties. Given $f\in Perf(X\times Y)$, it naturally induces a Fourier-Mukai transform
$\Gamma_{X\to Y}^{\mu}:Perf(X)\to Perf(Y),\quad\Gamma_{X\to Y}^{\mu}(\cdot)=\pi_{Y,*}(\pi_{X}^{*}(\cdot)\otimes\mu).$
Similarly, for any element $u\in\mathbb{H}^{*}(X\times Y,\mathbb{Q})$, one can define the cohomological correspondence
$\gamma_{X\to Y}^{u}:\mathbb{H}^{*}(X,\mathbb{Q})\to\mathbb{H}^{*}(Y,\mathbb{Q}),\quad\gamma_{X\to Y}^{u}(\cdot)=\pi_{Y*}(\pi_{X}^{*}(\cdot).\mu).$
Here $\mathbb{H}^{*}$ can be one of the Chow groups $CH^{*}$, the singular cohomology $H^{*}$, the Hodge cohomology $H^{*,*}$ or the Hochschild homology $HH_{*}$.
One may expect that two transforms can be naturally compared by the Chern chracter, i.e. we have a commutative diagram
(4.1) ${Perf(X)}$${Perf(Y)}$${\mathbb{H}^{*}(X,\mathbb{Q})}$${\mathbb{H}^{*}(Y,\mathbb{Q})}$$\scriptstyle{\Gamma_{X\to Y}^{\mu}}$$\scriptstyle{ch}$$\scriptstyle{ch}$$\scriptstyle{\gamma_{X\to Y}^{ch(\mu)}}$
where $ch$ is the Chern chracter. Unfortunately, 4.1 in general is not true due to the Grothendieck-Riemann-Roch theorem. Caldararu [4] found the correct way to modify the above diagram by generalizing the Mukai vectors, which we will recall in this section.
4.1. Fourier-Mukai transforms and cohomological correspondences
By the cohomological grading, we have a decomposition based on the odd or even degree
$H^{*}(X,\mathbb{Q})\cong H^{even}(X,\mathbb{Q})\oplus H^{odd}(X,\mathbb{Q})$
where $H^{even}(X,\mathbb{Q})$ is a commutative ring. By Section 2.4 of [4], there exists a unique formal series expansion such that
$\sqrt{1}=1,\quad\sqrt{uv}=\sqrt{u}\sqrt{v},\quad u=(\sqrt{u})^{2}$
for every smooth projective variety $X$ and any $u,v\in H^{even}(X,\mathbb{Q})$ (resp. $CH^{*}(X,\mathbb{Q})$ and $HH_{0}(X)$) with constant term $1$. We define $\hat{A}_{X}\in\mathbb{H}^{*}(X,\mathbb{Q})$ as
$\hat{A}_{X}:=td_{X}\cdot\sqrt{ch(K_{X})}$
where $K_{X}$ is the canonical line bundle of $X$ and $td_{X}$ is the Todd class of $X$.
Definition 4.1 (Definition 2.1 of [4]).
The directed Mukai vector of an element $\mu\in Perf(X\times Y)$ is defined by
$v(\mu,X\to Y):=ch(\mu)\cdot\sqrt{td_{X\times Y}}\cdot\sqrt[4]{ch(K_{Y})/ch(K_{X})}\in\mathbb{H}^{*}(X\times Y)$
By taking the first space to be a point we obtain the definition of the Mukai vector of an object $\alpha\in Perf(X)$ as
$v(\alpha)=ch(\alpha)\sqrt{\hat{A}_{X}}\in\mathbb{H}^{*}(X).$
Definition 4.2.
We define a map $(-)_{*}$ from the Fourier-Mukai transforms on $Perf(X\times Y)$ to cohomological correspondences $\mathbb{H}^{*}(X\times Y,\mathbb{Q})$ by
$\Gamma_{X\to Y*}^{\mu}:=\gamma_{X\to Y}^{v(\mu)}.$
Proposition 4.3 (Section 2.3 of [4]).
The map $(-)_{*}$ has the following properties:
  1. (1)
    Given Fourier-Mukai transforms $\mu:Perf(X)\to Perf(Y)$, $\nu:Perf(Y)\to Perf(Z)$, we have
    $\nu_{*}\circ\mu_{*}\cong\nu_{*}\circ\mu_{*};$
  2. (2)
    $id_{*}\cong id$;
  3. (3)
    Given a triangle of Fourier-Mukai transforms
    $f\to g\to h\to f[1],$
    we have $g_{*}=f_{*}+h_{*}$.
  4. (4)
    The following commutative diagram commutes:
    (4.2) ${Perf(X)}$${Perf(Y)}$${\mathbb{H}^{*}(X,\mathbb{Q})}$${\mathbb{H}^{*}(Y,\mathbb{Q})}$$\scriptstyle{\Gamma_{X\to Y}^{\mu}}$$\scriptstyle{v}$$\scriptstyle{v}$$\scriptstyle{\gamma_{X\to Y}^{v(\mu)}}$
4.2. Cohomological correspondences of derived Grassmannians
Now we come back to the setting of Section 3.2. Let $X$ be a smooth proper variety with a tor-amplitude $[0,1]$ perfect complex $\mathcal{E}$. We assume that for all $d\geq 0$, the derived Grassmannian
$Gr_{X}(\mathcal{E},d),\quad Gr_{X}(\mathcal{E}^{\vee}[1],d)$
are (classical) smooth varieties. We apply the Caldararu’s construction for $\Phi_{d_{+},d_{-}}$ and by 3.3 and 3.4 and have
Corollary 4.4.
The cohomological correspondences $\Phi_{d,d-1*},\Psi_{d,s-1*}^{R},\Phi_{d,d*},\Psi_{d,d*}^{L}$ satisfy the relations 2.2, if we denote
$E:=\Phi_{d,d-1*},\quad F:=\Psi_{d,d-1*}^{R},\quad K_{1}:=\Phi_{d,d*},\quad K_{-1}:=\Phi_{d,d*}^{L}.$
Now we consider the (co)homological degree of operators
$\Phi_{d,d-1*},\Psi_{d,s-1*}^{R},\Phi_{d,d*},\Psi_{d,d*}^{L}.$
For the Hochschild homology, their homological degree are all 00. But for the cohomology groups and Chow groups, their cohomological degree is not a constant and we need to make a modification. Recall that the incidence variety
$Inc_{X}(\mathcal{E},d_{+},d_{-})$
is quasi-smooth with virtual dimension $d_{0}=dim(X)+(1-d_{+})d_{+}-(d_{-}+1)d_{-}+d_{-}d_{+}$. By Ciocan-Fontanine and Kapranov [5], the derived structure induces a perfect obstruction theory on the classical scheme of $Inc(X,d_{+},d_{-})$ and thus induces a virtual fundamental class in the sense of Li-Tian [17] or Behrend-Fantechi [2]:
$Vir_{d_{+},d_{-}}:=[Inc_{X}(\mathcal{E},d_{+},d_{-})]^{vir}\in CH_{d_{0}}(Inc_{X}(\mathcal{E},d_{+},d_{-})).$
Now we define cohomological correspondences $\psi_{d,d-1},\psi_{d,d},\phi_{d,d-1},\phi_{d,d}$ such that
  1. (1)
    $\phi_{d,d}$ is induced by $Vir_{d,d}$, and $\psi_{d,d}$ is induced by $(-1)^{d}Vir_{d,d}$ but in a different direction.
  2. (2)
    both $\phi_{d,d-1}$ and $\psi_{d,d-1}$ are induced by $Vir_{d,d-1}$ but in two different directions.
Proposition 4.5.
The operators $\phi_{d,d-1}$ and $\psi_{d,d-1}$ presevered the cohomological degree and the operators $\phi_{d,d}$ and $\psi_{d,d}$ shifted the Chow group (resp. cohomological, Hodge) degree by $d$ (resp. $2d$, $(d,d)$) and $-d$ (resp. $-2d$, $(-d,-d)$). Moreover, the cohomological correspondences $\phi_{d,d-1},\psi_{d,s-1},\phi_{d,d},\psi_{d,d}$ also satisfy the relations 2.2 if we denote
$E:=\phi_{d,d-1},\quad F:=\psi_{d,d-1},\quad K_{1}:=\phi_{d,d},\quad K_{-1}:=\phi_{d,d}.$
Proof.
First we notice that all the constant terms of roots of Todd class and Chern classes of line bundles are $1$. By the virtual Grothendieck-Riemann-Roch theorem (i.e. Theorem 3.23 and Corollary 3.25 of [14]), we also have
$ch(\mathcal{O}_{Inc_{X}(\mathcal{E},d_{+},d_{-})})=Vir_{d_{+},d_{-}}+C_{d_{+},d_{-}}$
where $C_{d_{+},d_{-}}$ have higher degrees than $Vir_{d_{+},d_{-}}$. Thus we have $\Phi_{d,d*}-\phi_{d,d}$ has higher cohomological degree than $\phi_{d,d}$. Similar computations also holds for $\Phi_{d,d-1*}-\phi_{d,d-1},\Psi_{d,d*}^{L}-\psi_{d,d}$ and $\Psi_{d,d*}-\psi_{d,d-1}$. Thus by checking the cohomological grading, the relations 2.2 for $\phi_{d,d-1},\psi_{d,s-1},\phi_{d,d},\psi_{d,d}$ follows from it for $\Phi_{d,d-1*},\Psi_{d,d-1*}^{R},\Phi_{d,d*},\Psi_{d,d*}^{L}$. ∎
In a companion paper [29], we will prove that
Proposition 4.6.
The incidence variety
$Inc_{X}(\mathcal{E},d_{+},d_{-})$
is classical and
$Vir_{d_{+},d_{-}}:=[Inc_{X}(\mathcal{E},d_{+},d_{-})]\in CH_{d_{0}}(Inc_{X}(\mathcal{E},d_{+},d_{-})).$
5. Perverse Coherent Sheaves and their moduli spaces
Let $S$ be a smooth surface with a fixed ample line bundle $H$. We denote by
$p:\hat{S}\to S$
the blow up of $S$ at a closed point $o\in S$. We denote $C:=p^{-1}(o)$ as the exceptional (rational) curve. We denote
$\mathcal{O}_{C}(m):=\mathcal{O}(-mC)|_{C}$
which is the unique degree $m$ line bundle on the rational curve $C$ for any $m\in\mathbb{Z}$. Given an ample line bundle $H$ on $S$, Nakajima-Yoshioka defined the perverse coherent sheaf as
Definition 5.1.
A stable perverse coherent sheaf $E$ on $\hat{S}$ with respect to $H$ is a coherent sheaf such that
$Hom(E,\mathcal{O}_{C}(-1))=0,\quad Hom(\mathcal{O}_{C},E)=0$
and $p_{*}E$ is Gieseker stable with respect to $H$.
Given a coherent sheaf $E$ on $\hat{S}$ and an integer $m$, we define $E$ to be $m$-stable if $E\otimes\mathcal{O}(-mc)$ is a perverse stable sheaf.
Remark 5.2.
When the rank of $E$ is $1$, $p_{*}E$ is stable with respect to $H$ if and only if $p_{*}E$ is torsion-free, and hence independent with the choice of $H$.
By the work of Nakajima-Yoshioka [23, 22, 24], the stable perverse coherent sheaf builds a bridge between the Hilbert scheme of points on $S$ and $\hat{S}$ (and more generally the moduli space of stable sheaves), which we recall in this section.
5.1. Moduli space of perverse coherent sheaves
We recall the decomposition of cohomology groups:
$H^{*}(\hat{S},\mathbb{Q})=p^{*}(H^{*}(\hat{S},\mathbb{Q}))\oplus\mathbb{Q}[C]$
where $[C]\in H^{1,1}(\hat{S})\cap H^{2}(\hat{S},\mathbb{Q})$ and $[C]^{2}=-1$.
Definition 5.3.
Given Chern characters $\hat{c}\in H^{*}(\hat{S},\mathbb{Q})$, we denote $M^{m}(\hat{c})$ as the moduli space of $m$-stable coherent sheaves on $\hat{S}$ with Chern character $\hat{c}$. Particularly, given integers $l,n$, we denote
$M^{m}(l,n):=M^{m}([\hat{S}]-l[C]+(-n-\frac{l^{2}}{2})[pt]),$
i.e. it consists of $m$-stable perverse coherent sheaves $E$ such that
$rank(E)=1,\quad c_{1}(E)=-l[C],\quad c_{2}(E)=n$
By tensoring $\mathcal{O}(-mC)$, we have a canonical isomorphism
$M^{m}(\hat{c})\cong M^{0}(\hat{c}e^{-m[C]}),\quad M^{m}(l,n)\cong M^{0}(l+m,n)$
Lemma 5.4 (Lemma 3.3 of [15] and Proposition 3.37 of [22] ).
For any integers $l,n$, the variety $M^{0}(l,n)$ is either empty or a smooth projective variety of $2n$ dimension. Particularly, fixing $n$, we have
$M^{0}(0,n)\cong S^{[n]},\quad M^{0}(l,n)\cong\hat{S}^{[n]}\text{ if }l\gg 0.$
5.2. Moduli spaces as derived Grassmannians
We first recall cohomological properties of stable coherent sheaves
Proposition 5.5 (Proposition 3.1 of [16] and Lemma 3.4 of [22]).
For a stable perverse coherent sheaf $E$ on $\hat{S}$, we have
$R^{i}p_{*}E(aC)=0,\quad a=0,1,\quad i\geq 1$
and moreover
$Rp_{*}E(aC)=p_{*}E(aC),\quad a=0,1$
are all stable coherent sheaves with respect to $H$.
For any stable perverse coherent sheaf $E$ such that
$rank(E)=1,\quad c_{1}(E)=-l[C],\quad c_{2}(E)=n,$
by Proposition 5.5
$p_{*}E,\text{ and }p_{*}E(C)$
are all stable and by the Grothendieck-Riemann-Roch theorem we have
$rank(p_{*}E)=1,\quad c_{1}(p_{*}E)=0,\quad c_{2}(p_{*}E)=(n+\frac{l+l^{2}}{2}).$
$rank(p_{*}E(C))=1,\quad c_{1}(p_{*}E(C))=0,\quad c_{2}(p_{*}E(C))=(n+\frac{-l+l^{2}}{2}).$
It induces morphisms of moduli spaces
$\zeta:M^{0}(l,n)\to S^{[n+\frac{l+l^{2}}{2}]},\quad E\to p_{*}(E)$
$\eta:M^{0}(l,n)\to S^{[n+\frac{-l+l^{2}}{2}]},\quad E\to p_{*}(E(C)).$
The morphisms $\zeta$ and $\eta$ are actually projection morphisms of derived Grassmannians of the universal ideal sheaves over Hilbert schemes of points on $S$: let
$\mathcal{I}_{m}\in Coh(S^{[m]}\times S)$
be the universal ideal sheaf of $S^{[m]}$ and we denote
$\mathcal{I}_{m,o}:=\mathcal{I}_{m}|_{S^{[m]}\times o},\quad(\text{short for }\mathcal{I}_{o}).$
By a folklore lemma (like Proposition 2.14 of [25]), $\mathcal{I}_{o}$ is of tor amplitude $[0,1]$.
Theorem 5.6 (Theorem 4.1 of [22]).
We have a canonical isomorphism
$M^{0}(l,n)\cong Gr_{S^{[n+\frac{-l+l^{2}}{2}]}}(\mathcal{I}_{o},l),\quad M^{0}(l,n)\cong Gr_{S^{[n+\frac{l+l^{2}}{2}]}}(\mathcal{I}_{o}^{\vee}[1],l),$
where the projection morphisms are $\eta$ and $\zeta$ exactly.
6. The admissible $\mathcal{E}$-representation on perverse coherent sheaves
Given a smooth projective surface $S$ and a closed point $o\in S$ and a cohomology theory $\mathbb{H}$, we consider
$\mathbb{H}^{S,o}_{l}:=\bigoplus_{n\in\mathbb{Z}}\mathbb{H}^{*}(M^{0}(l,n),\mathbb{Q}),\quad\mathbb{H}^{S,o}:=\bigoplus_{l\in\mathbb{Z}}\mathbb{H}^{S,o}_{l}.$
When $l=0$, we have
$\mathbb{H}_{0}^{S,o}\cong\bigoplus_{n\in\mathbb{Z}}\mathbb{H}^{*}(S^{[n]},\mathbb{Q}).$
Other than the grading by the first Chern characters (or equivalently $m$-stability), we also have two gradings: the $q$-degree by associating $\mathbb{H}^{*}(M^{0}(l,n),\mathbb{Q})$ with degree $n$, the $t$-degree (or bi-degree) which is the cohomology degree. In this section, we prove that
Theorem 6.1 (Theorem 1.6).
There is a $\mathcal{E}$-representation on $\mathbb{H}_{0}^{S,o}$ such that
  1. (1)
    the representation is graded admissible;
  2. (2)
    the $q$-degree of $E,F$ are preserved, and the $q$-degree of $K_{\pm 1}$ are ±$\pm l$ on $\mathbb{H}^{S,o}_{l}$
  3. (3)
    the cohomology degree of $E,F$ are preserved, and the cohomology degree (with respect to Hochschild homology, Hodge cohomolgy and Chow group) of $K_{\pm 1}$ are $0,\pm(l,l),\pm l$ on $\mathbb{H}^{S,o}_{l}$.
  4. (4)
    the stable limit
    $\mathbb{H}_{\infty}^{S,o}\cong\bigoplus_{n\in\mathbb{Z}}\mathbb{H}^{*}(\hat{S}^{[n]},\mathbb{Q}).$
  5. (5)
    the operators are given by geometric correspondences.
Once we have the above graded admissible representation, by the categorical Boson-Fermion correspondence theorem (i.e. Theorem 2.21), we have
Theorem 6.2 (Theorem 1.3).
We have an identification
$\mathbb{H}^{s,o}\cong\mathbb{H}_{0}^{S,o}\otimes_{\mathbb{Q}}F.$
and the action of the infinite dimensional Clifford algebra on $F$ is given by geometric correspondences. The cohomological degree (with respect to Hoschild homology, Hodge cohomolgy and Chow group) of a monomial in $F$:
$x_{i_{1}}\wedge\cdots x_{i_{n}}$
is $0,(x_{i_{1}}+\cdots+x_{i_{n}}-\frac{n^{2}+n}{2},x_{i_{1}}+\cdots+x_{i_{n}}-\frac{n^{2}+n}{2}),x_{i_{1}}+\cdots+x_{i_{n}}-\frac{n^{2}+n}{2}$ respectively and the $q$-degree is equal to the $t$-degree of the Chow group.
Theorem 6.3 (Theorem 1.1).
Let $\mathbb{H}^{*}$ be the functor of singular cohomology groups, Hochschild homology groups, or the Chow groups. We have an identification
(6.1) $\mathbb{H}_{\infty}^{S,o}\cong\mathbb{H}_{0}^{S,o}\otimes_{\mathbb{Q}}B.$
In particular, the Heisenberg algebra action on $B$ is induced by geometric correspondences. For a monomial in the Bosonic Fock space
$y_{1}^{i_{1}}y_{2}^{i_{2}}\cdots y_{n}^{i_{n}}$
its cohomological degree with respect to Hoschild homology, Hodge cohomology and Chow group is
$0,\quad\sum_{j=1}^{n}ji_{j}(1,1),\quad\sum_{j=1}^{n}ji_{j}.$
and the $q$-degree is same as the $t$-degree of the Chow group.
6.1. The geometric correspondences
We recall that over $S^{[n+\frac{-l+l^{2}}{2}]}$, by Theorem 5.6 the derived Grassmannian of $\mathcal{I}_{o}$ induces
$Gr(\mathcal{I}_{o}^{\vee}[1],l)\cong M^{0}(l,n-l),\quad Gr(\mathcal{I}_{o}^{\vee}[1],l-1)\cong M^{0}(l-1,n),\quad Gr(\mathcal{I}_{o},l)\cong M^{0}(l,n).$
Thus the operators
$\psi_{l,l-1}\quad\phi_{l,l-1}\quad\psi_{l,l}\quad\phi_{l,l}$
acts on
$\psi_{l,l-1}:\bigoplus_{n\geq 0}\mathbb{H}^{*}(M^{0}(l,n),\mathbb{Q})\to\bigoplus_{n\geq 0}\mathbb{H}^{*}(M^{0}(l-1,n),\mathbb{Q}),$
$\phi_{l-1,l}:\bigoplus_{n\geq 0}\mathbb{H}^{*}(M^{0}(l-1,n),\mathbb{Q})\to\bigoplus_{n\geq 0}\mathbb{H}^{*}(M^{0}(l,n),\mathbb{Q}),$
$\psi_{l,l}:\bigoplus_{n\geq 0}\mathbb{H}^{*}(M^{0}(l,n),\mathbb{Q})\to\bigoplus_{n\geq 0}\mathbb{H}^{*}(M^{0}(l,n-l),\mathbb{Q}),$
$\phi_{l,l-1}:\bigoplus_{n\geq 0}\mathbb{H}^{*}(M^{0}(l,n-l),\mathbb{Q})\to\bigoplus_{n\geq 0}\mathbb{H}^{*}(M^{0}(l,n),\mathbb{Q}).$
We define the operators
(6.2) $E:=\bigoplus_{l\geq 0}\phi_{l-1,l},\quad F:=\bigoplus_{l\geq 0}\psi_{l-1,l},$
$K_{1}:=\bigoplus_{l\geq 0}\phi_{l,l},\quad K_{-1}:=\bigoplus_{l\geq 0}\psi_{l,l}.$
Then they all acts on $\mathbb{H}_{l,n}^{S,o}$ (for the Hochschild homology, we replace it by $\Psi_{*,*}$ and $\Phi_{*,*}$ respectively).
Proof of Theorem 6.1.
By Proposition 4.5, operators in 6.2 forms an $\mathcal{E}$-representation by geometric correspondences and the $q,t$-degree matches the Theorem 6.1. The admissibility follows from the fact that
$M(l,n)=\emptyset$
when $n<0$. ∎
Appendix A Some algebra computations
In this section, we give the proof of Proposition 2.2, Proposition 2.16, Proposition 2.17 and Proposition 2.19, which involve explicit computations.
A.1. The proof of Proposition 2.2
We recall the definition of the $P_{i},Q_{i}$ for $i>0$ in the algebra $\mathcal{E}$:
Definition A.1.
We define elements $P_{i},Q_{i}$ in $\mathcal{E}$ for any $i>0$ such that $P_{1}=EK_{-1},Q_{1}=K_{1}F$ and
(A.1) $P_{i}=K_{1}P_{i-1}K_{-1}-EP_{i-1}F,\quad Q_{i}=K_{1}Q_{i-1}K_{-1}-EQ_{i-1}F,\quad\text{ if }i>1.$
Proof of Proposition 2.2.
If we denote
$P(z):=\sum_{i=1}^{\infty}P_{i}z^{i},\quad Q(z):=\sum_{i=1}^{\infty}Q_{i}z^{i}$
Then $P(z),Q(z)$ are decided by $P_{1},Q_{1}$ and the equation
(A.2) $P(z)=z(P_{1}+K_{1}P(z)K_{-1}-EP(z)F)$
(A.3) $Q(z)=z(Q_{1}+K_{1}Q(z)K_{-1}-EQ(z)F).$
We need to prove that
$P(z)P(w)+P(w)P(z)=0,$
$Q(z)Q(w)+Q(w)Q(z)=0,$
$P(z)Q(w)+Q(w)P(z)=\frac{zw}{1-zw}.$
By A.2, we have
$P(z)K_{1}=z(E+K_{1}P(z))\quad K_{-1}P(z)=zP(z)K_{-1},$
$P(z)E=-zEP(z),\quad FP(z)=z(K_{-1}-P(z)F),$
$Q(z)K_{1}=zK_{1}Q(z),\quad K_{-1}Q(z)=z(F+K_{-1}Q(z)),$
$Q(z)E=z(K_{1}-EQ(z)),\quad FQ(z)=-zQ(z)F.$
and similar equations for $Q(z)$. Moreover, we have
$P_{1}P(z)=EK_{-1}P(z)=zEP(z)K_{-1}=-P(z)P_{1},$
$P_{1}Q(z)=EK_{-1}Q(z)=zE(F+Q(z)K_{-1})$
$=-Q(z)P_{1}+z(EF+K_{1}K_{1})=-Q(z)P_{1}+z.$
and similarly
$Q_{1}P(z)=-P(z)Q_{1}+z,\quad Q_{1}Q(z)=-Q(z)Q_{1}.$
Hence we have
(A.4) $P(z)P(w)=zw(P_{1}+K_{1}P(z)K_{-1}-EP(z)F)(P_{1}+K_{1}P(w)K_{-1}-EP(w)F)$
$=zP_{1}P(w)+wP(z)P_{1}-zwP_{1}^{2}+zw(K_{-1}P(z)P(w)K_{1}+EP(z)P(w)F).$
We notice that $P_{1}^{2}=EK_{-1}EK_{1}=0$. Thus let $\mathfrak{P}=P(z)P(w)+P(w)P(z)$. We have
$\mathfrak{P}=z(P_{1}P(w)+P(w)P(z))+w(P_{1}P(z)+P(z)P_{1})+zw(K_{-1}\mathfrak{P}K_{1}+F\mathfrak{P}E)$
$=zw(K_{-1}\mathfrak{P}K_{1}+F\mathfrak{P}E)$
By the boundary condition, the only solution of
$\mathfrak{P}=zw(K_{-1}\mathfrak{P}K_{1}+F\mathfrak{P}E)$
is 00 and thus
$P(z)P(w)=P(w)P(z).$
Similarly we have
$Q(z)Q(w)=Q(w)Q(z).$
Similar to A.4, we have
$P(z)Q(w)=zP_{1}Q(w)+wP(z)Q_{1}-zwP_{1}Q_{1}+zw(K_{-1}P(z)Q(w)K_{1}+EP_{-1}(z)P(w)F),$
$Q(w)P(z)=wQ_{1}P(z)+zQ(w)P_{1}-zwQ_{1}P_{1}+zw(K_{-1}Q(z)P(w)K_{1}+EQ(z)P(w)F).$
We notice that
$P_{1}Q_{1}=EF,\quad Q_{1}P_{1}=K_{-1}K$
and thus $P_{1}Q_{1}+Q_{1}P_{1}=1$. Hence let $\mathfrak{Q}:=P(z)Q(w)+Q(w)P(z)$ and we have
$\mathfrak{Q}=zw(1+K_{-1}\mathfrak{Q}K_{1}+E\mathfrak{Q}F)$
and hence
$(\mathfrak{Q}-\frac{zw}{1-zw})=zw(K_{-1}(\mathfrak{Q}-\frac{zw}{1-zw})K_{1}+E(\mathfrak{Q}-\frac{zw}{1-zw})F)$
and hence
$P(z)Q(w)+Q(w)P(z)=\frac{zw}{1-zw}.$
A.2. The proof of Proposition 2.16
Now let $W$ be an graded adimissible $\mathcal{C}$-module. To prove Proposition 2.16, we first prove the following lemmas
Lemma A.2.
Let $\beta$ be a degree 00 operator on $W$, such that $\beta|_{W_{0}}=0$ and $[Q(z),\beta]=0$, i.e.
$[Q_{l},\beta]=0$
for any $l>0$, then $\beta=0$.
Proof.
We prove that $\beta|_{W_{i}}=0$ by the induction on $i$. Assuming that $\beta|_{W_{i}}=0$, then on $W_{i+1}$,
$Q_{l}\beta|_{W_{i+1}}=0$
for all $l>0$ and hence $\beta=0$ by the admissible property of $W$. ∎
Corollary A.3.
Let $\alpha$ and $\beta$ be two degree 00 operators on $W$ such that $\beta|_{W_{0}}=\alpha|_{W_{0}}$ and
$[Q_{l},\beta]=[Q_{l},\alpha]$
for all $l>0$. Then $\alpha=\beta$.
Recalling
$\Lambda:=\Gamma_{0}(id).$
Lemma A.4.
We have $\Lambda|_{W_{i}}=i\cdot id$.
Proof.
We notice that
$Q_{l}\Lambda=-\sum_{i=1}^{\infty}P_{i}Q_{l}Q_{i}+Q_{l}=\Lambda Q_{l}+Q_{l}$
and thus $\Lambda$ and $\oplus_{i\geq 0}i\cdot id$ all satisfy the equation
$[Q_{l},*]=Q_{l}$
for all $l>0$. Hence they are the same by Corollary A.3.
Lemma A.5.
We have
$Q(z)K_{1}=zK_{1}Q(z),\quad K_{-1}P(z)=zP(z)K_{-1},$
i.e. if $j>0$
$Q_{j}K_{1}=K_{1}Q_{j-1},\quad K_{-1}P_{j}=P_{j-1}K_{-1},$
where we denote $P_{0}=Q_{0}=0$.
We also have
$P(z)K_{1}=zK_{1}P(z)+P_{1}K_{1},\quad K_{-1}Q(z)=zK_{-1}Q(z)+K_{-1}Q_{1},$
i.e. if $j>1$
$P_{j}K_{1}=K_{1}P_{j-1},\quad K_{-1}Q_{j}=Q_{j-1}K_{-1}.$
Proof.
We only prove $Q_{j}K_{1}=K_{1}Q_{j-1}$ and other computations follow from the same method. We notice that
$(\Lambda+id)Q_{j}K_{1}=Q_{j}\Lambda K_{1}=\sum_{l=1}^{\infty}Q_{j}P_{l+1}K_{1}Q_{l}=-\sum_{j=1}^{\infty}P_{l+1}(Q_{j}K_{1})Q_{l}+K_{1}Q_{j-1}$
$\Lambda K_{1}Q_{j-1}=\sum_{l=1}^{\infty}P_{l+1}K_{1}Q_{l}Q_{j-1}=-\sum_{j=1}^{\infty}P_{l+1}(K_{1}Q_{j-1})Q_{j-1}$
Thus $\alpha:=Q_{j}K_{1}-K_{1}Q_{j}$ satisfy the equation that
(A.5) $(\Lambda+id)\alpha=-\Gamma_{1}(\alpha).$
By induction, the solutions of A.5 is decided by $\alpha|_{W_{0}}$, and hence $\alpha=0$ follows from $\alpha|_{W_{0}}=0$. ∎
Corollary A.6.
For any $j>0$ (resp. $j>1$), we have
$Q_{j}K_{-1}K_{1}=K_{-1}K_{1}Q_{j},\quad Q_{j}K_{1}K_{-1}=K_{1}K_{-1}Q_{j},$
i.e.
(A.6) $[Q(z),K_{-1}K_{1}]=0,\quad[\frac{d}{dz}Q(z),K_{1}K_{-1}]=0.$
Proof.
It directly follows from Lemma A.5. ∎
Corollary A.7.
On $W$ we have
$K_{1}K_{-1}=Q_{1}P_{1},\quad K_{-1}K_{1}=id.$
Proof.
The second formula follows from A.6 and Lemma A.2. For the first formula, let $\beta:=K_{1}K_{-1}-Q_{1}P_{1}$. We have
$\beta|_{W_{0}}=0,\quad Q_{1}\beta=0,\quad[Q_{l},\beta]=0$
for $l>1$. Then by induction and the similar method with Lemma A.2, we also have $\beta=0$. ∎
Proof of Proposition 2.16.
The formula $K_{-1}K_{1}=id$ follows from Corollary A.7. We have
$FE=K_{-1}P_{1}Q_{1}K_{1}=K_{-1}K_{1}K_{-1}K_{1}=id$
by Corollary A.7. Also by Corollary A.7,
$K_{1}K_{-1}+EF=Q_{1}P_{1}+P_{1}Q_{1}P_{1}Q_{1}=Q_{1}P_{1}+P_{1}Q_{1}-P_{1}^{2}Q_{1}^{2}=id.$
Finally we have
$FK_{1}=K_{-1}P_{1}K_{1}=0,\quad K_{-1}E=K_{-1}Q_{1}K_{1}=0$
by Lemma A.5. ∎
A.3. The proof of Proposition 2.17
Let $H_{\pm k}$, $k\in\mathbb{Z}-\{0\}$ be the generators of the infinite dimensional algebra. By induction we have
Lemma A.8.
For any integer $i\geq 0$, we have
(A.7) $H_{-k}H_{k}^{i}=H_{k}^{i}H_{-k}-ik,\quad H_{-k}^{i}H_{k}=H_{k}H_{-k}^{i}-ik.$
Let $W_{\infty}$ be an admissible $\mathcal{H}$-representation, with the filtration
$W_{i}:=\bigcap_{l>i}ker(H_{l}).$
and the operators for every $k>0$
$\mathcal{L}_{k}:=-\sum_{i=1}^{\infty}\frac{1}{k^{i}i!}H_{k}^{i-1}H_{-k}^{i}.$
Lemma A.9.
We have
(A.8) $\mathcal{L}_{k}H_{k}=id,\quad H_{-k}(1-H_{k}\mathcal{L}_{k})=0.$
Proof.
By A.7 we have
$\mathcal{L}_{k}H_{k}=-\sum_{i\geq 1}\frac{1}{k^{i}i!}H_{k}^{i-1}H_{-k}^{i}H_{k}=\sum_{i\geq 1}(\frac{1}{k^{i-1}(i-1)!}H_{k}^{i-1}H_{-k}^{i-1}-\frac{1}{k^{i}i!}H_{k}^{i}H_{-k}^{i})=1.$
and
$H_{-k}(1-H_{k}\mathcal{L}_{k})=H_{-k}(Id+\sum_{i=1}^{\infty}\frac{1}{k^{i}i!}H_{k}^{i}H_{-k}^{i})$
$=H_{-k}+\sum_{i=1}^{\infty}(\frac{1}{k^{i}i!}H_{k}^{i}H_{-k}^{i+1}-\frac{1}{k^{i-1}(i-1)!}H_{k}^{i-1}H_{-k}^{i})=H_{-k}-H_{-k}=0.$
By A.8, we have
$(Id-H_{k}\mathcal{L}_{k})W_{k}\subset W_{k-1}.$
Now we consider
$W:=\bigoplus_{i\geq 0}W_{i}.$
and define degree 00 operators
$K_{1}=\bigoplus_{i\geq 0}H_{k},\quad K_{-1}=\bigoplus_{i\geq 0}\mathcal{L}_{k}$
where $H_{0}=\mathcal{L}_{0}:=id$.
The inclusion of $W_{k}\subset W_{k+1}$ naturally induces a degree $1$ operator on $W$:
$E:=\bigoplus_{k\geq 0}id.$
By A.8, the following degree $-1$ operator is well defined
$F:=\bigoplus_{k\geq 0}(Id-H_{k}\mathcal{L}_{k})$
Proof of Proposition 2.17.
We need to prove that under the above construction, the vector space $W$ forms an admissible graded $\mathcal{E}$-algebra representation.
First we notice that for any $k>0$, $\mathcal{L}_{k}|_{W_{k-1}}=0$ and thus $K_{-1}E=0$ and $FE=id$. The formula $K_{1}K_{-1}+EF=0$ follows from the definition of $K_{\pm 1}$ and $E,F$.
On the other hand, by A.8 we have
$K_{-1}K_{1}=id,\quad FK_{1}=(1-K_{1}K_{-1})K_{1}=0.$
Thus $W$ is a graded $\mathcal{E}$-representation and the admissibility of $W$ follows from the admissibility of $\mathcal{W}_{\infty}$. ∎
A.4. The proof of Proposition 2.19
Let $W$ be an admissible graded $\mathcal{E}$-representation. As the limit of 2.3, on $W_{l}$ ($l>0$) we have
(A.9) $\sum_{l=0}^{\infty}E_{l}F_{l}=id.$
Lemma A.10.
For any $i>0$, the operators $H_{i,i}$ and $H_{i,-i}$ satisfy the relation
$H_{i,i}H_{i-i}-H_{i,-i}H_{i,i}=i\cdot id.$
Proof.
We have
$H_{i,i}H_{i,-i}=-iK_{1}\sum_{l=0}^{\infty}K_{1}^{l}K_{-1}^{l+1}=i-\sum_{l=0}^{\infty}(K_{1})^{l}K_{-1}^{l}$
$=i-\sum_{l=0}^{\infty}(K_{1})^{l}K_{-1}^{l+1}K_{1}=i+H_{i,-i}H_{i,i}.$
Lemma A.11.
For any $i>0$ and $|j|<i$, we have
$H_{i,j}H_{i,i}=H_{i,i}H_{i,j},\quad H_{i,j}H_{i,-i}=H_{i,-i}H_{i,j}$
Proof.
First we prove that $H_{i,j}H_{i,i}=H_{i,i}H_{i,j}$. It follows from
$H_{i,j}H_{i,i}=\sum_{l=0}^{\infty}E_{l}H_{i-1,j}F_{l}K_{1}=\sum_{l=0}^{\infty}E_{l+1}H_{i-1,j}F_{l}=H_{i,i}H_{i,j}$
by Lemma 2.1. Then we prove that $H_{i,j}H_{i,-i}=H_{i,-i}H_{i,j}$. It follows from
$H_{i,j}H_{i,-i}=-i\sum_{l,m\geq 0}E_{l}H_{i-1,j}F_{l}K_{1}^{m}K_{-1}^{m+1}=-i\sum_{l\geq m\geq 0}E_{l}H_{i-1,j}F_{l+1},$
$H_{i,-i}H_{i,j}=-i\sum_{l,m\geq 0}K_{1}^{m}K_{-1}^{m+1}E_{l}H_{i-1,j}F_{l}=-i\sum_{l>m\geq 0}E_{l-1}H_{i-1,j}$
$=-i\sum_{l\geq m\geq 0}E_{l}H_{i-1,j}F_{l+1}=H_{i,j}H_{i,-i}$
by Lemma 2.1. ∎
Proposition A.12.
On $W_{i}$, the operators $H_{i,j}$ where $j\neq 0$ and $-i\leq j\leq i$ satisfy the relation
(A.10) $H_{i,j}H_{i,k}-H_{i,k}H_{i,j}=j\delta_{j,-k}id.$
Proof.
We assume that $|j|\leq|k|$. We prove it by induction on $i$. It is trivial when $i=0$. Assuming it is true for $i-1$, then if $k=\pm i$, A.10 follows from Lemma A.10 and Lemma A.11. Otherwise, by Lemma 2.1 we have
$H_{i,j}H_{i,k}=\sum_{l,m\geq 0}E_{l}H_{i-1,j}F_{l}E_{m}H_{i-1,k}F_{m}=E_{l}H_{i-1,j}H_{i-1,k}F_{l}.$
Thus we have
$[H_{i,j},H_{i,k}]=\sum_{l=0}^{\infty}E_{l}[H_{i-1,j},H_{i-1,k}]F_{l}=j\delta_{j,-k}\sum_{l=0}^{\infty}E_{l}F_{l}=j\delta_{j,-k}id$
by A.9. Here [,][-,-][ - , - ] denotes the Lie bracket. ∎
Proof of Proposition 2.19 .
It is the limit of Proposition A.12 when $i\to\infty$. ∎