### v3

# Hilbert schemes on blowing ups and the free Boson

Abstract.
For different cohomology theories (including Hochschild homology, Hodge cohomology, Chow groups, and Grothendieck groups of coherent sheaves), we identify the cohomology of the moduli space of rank $1$ perverse coherent sheaves on the blow-up of a surface by Nakajima-Yoshioka [22] as the tensor product of the cohomology of Hilbert schemes with a Fermionic Fock space. As the stable limit, we identify the cohomology of Hilbert schemes of the blow-up of a surface as the tensor product of the cohomology of Hilbert schemes on the surface with a Bosonic Fock space. The actions of the infinite-dimensional Clifford algebra and Heisenberg algebra are all given by geometric correspondences.
1. Introduction
1.1. Motivation and the Main Theorem
Let $S$ be a smooth projective surface over $\mathbb{C}$. The famous Göttsche’s formula computed the Poincaré polynomials of Hilbert schemes of points on $S$ [11]:
(1.1) $\sum_{n=0}^{\infty}q^{n}P_{t}(S^{[n]})=\prod_{m=1}^{\infty}\frac{(1+t^{2m-1}q^{m})^{b_{1}(S)}(1+t^{2m+1}q^{m})^{b_{3}(S)}}{(1-t^{2m-2}q^{m})^{b_{0}(S)}(1-t^{2m}q^{m})^{b_{2}(S)}(1-t^{2m+2}q^{m})^{b_{4}(S)}},$
where $b_{i}(S)$ is the $i$-th Betti number.
  1. (1)
    By letting $t=1$, the formula above is the character of a super-Fock space. Nakajima [21] and Grojnowski [10] realized the cohomology group of the Hilbert scheme as the Fock space of a super-Heisenberg algebra through geometric correspondences.
    By replacing the cohomology with the Chow group, the Heisenberg (and moreover, Virasoro and $W_{1+\infty}$) algebra action was constructed by Maulik-Negut [20]. However, as the Chow group of a surface can be infinite-dimensional (like $K3$ surfaces), the Chow groups of all $S^{n}$ are involved to determine the Chow group of Hilbert schemes, and the representation theory is also more complicated (for example, $sp_{2\infty}$ will naturally appear).
  2. (2)
    Let $\hat{S}$ be the blow-up of $S$ at a point $o$. By Göttsche’s formula, we have
    $\sum_{n=0}^{\infty}q^{n}P_{t}(\hat{S}^{[n]})=\prod_{m=1}^{\infty}\frac{1}{1-t^{2m}q^{m}}\sum_{n=0}^{\infty}q^{n}P_{t}(S^{[n]}),$
    where the ratio
    $\prod_{m=1}^{\infty}\frac{1}{1-t^{2m}q^{m}}$
    is the character of a Bosonic Fock space. Thus, it suggests we have a canonical isomorphism
    (1.2) $\bigoplus_{n=0}^{\infty}H^{*}(\hat{S}^{[n]},\mathbb{Q})\cong\bigoplus_{n=0}^{\infty}H^{*}(S^{[n]},\mathbb{Q})\otimes_{\mathbb{Q}}B,$
    where $B$ is the Bosonic Fock space $\mathbb{Q}[y_{1},y_{2},\cdots]$. The formula (1.2) does not depend on the choice of the surface $S$ and the point $o$, and might also hold if we consider the generalized cohomology theory like Hochschild homology, Chow groups and Grothendieck group of coherent sheaves.
The purpose of this paper is to realize the expectation (1.2) by proving that:
Theorem 1.1 (Theorem 6.3).
Let $\mathbb{H}^{*}$ be the functor of Hodge cohomology groups, Hochschild homology groups, Chow groups, or Grothendieck groups of coherent sheaves. We have an identification
(1.3) $\bigoplus_{n=0}^{\infty}\mathbb{H}^{*}(\hat{S}^{[n]},\mathbb{Q})\cong\bigoplus_{n=0}^{\infty}\mathbb{H}^{*}(S^{[n]},\mathbb{Q})\otimes_{\mathbb{Q}}B,$
where the Heisenberg algebra action on $B$ is induced by geometric correspondences.
Remark 1.2.
More details of our geometric correspondences will be given in Section 1.2 to Section 1.4. At the current stage, we do not know the precise relation between our Heisenberg algebra action and that of Nakajima [21] and Grojnowski [10] in cohomology groups or Maulik-Neguţ [20] in Chow groups. We expect the study of the affine super-Yangian will shed some light on this problem. See Section 1.5 for more discussions.
1.2. Stable Limits and Perverse Coherent Sheaves
Theorem 1.1 follows from regarding both the Bosonic Fock space $B$ and $\mathbb{H}^{*}(\hat{S}^{[n]},\mathbb{Q})$ as a stable limit:
  1. (1)
    We consider the Fermionic Fock space $F$ which is the exterior algebra of
    $V=\bigoplus_{i=1}^{\infty}\mathbb{Q}x_{i}.$
    The degree $1$ operator $K_{1}$
    $K_{1}:F_{i}\to F_{i+1},\quad x_{j_{1}}\wedge\cdots\wedge x_{j_{i}}\to x_{0}\wedge x_{j_{1}+1}\wedge\cdots\wedge x_{j_{i}+1}$
    is injective. Moreover, by direct computation, the stable limit
    $F_{\infty}:=\lim_{i\to\infty}F_{i}\cong B.$
  2. (2)
    Nakajima-Yoshioka [22] considered the moduli space of $l$-stable perverse coherent sheaves $M^{l}(c_{n})$ on $\hat{S}$, where $l\in\mathbb{Z}$ and
    $c_{n}=[\hat{S}]+n[o]$
    is the Chern character. It has the property that
    $M^{0}(c_{n})\cong S^{[n]},\quad M^{l}(c_{n})\cong\hat{S}^{[n]}\text{ if }l\gg 0.$
    Hence $\hat{S}^{[n]}$ can be regarded as the stable limit of $M^{l}(c_{n})$, $l\to\infty$.
Theorem 1.1 follows from the representation theory on $M^{l}(c_{n})$:
Theorem 1.3 (Theorem 6.2).
We have an identification
(1.4) $\bigoplus_{l=0}^{\infty}\bigoplus_{n=0}^{\infty}\mathbb{H}^{*}(M^{l}(c_{n}),\mathbb{Q})\cong\bigoplus_{n=0}^{\infty}\mathbb{H}^{*}(S^{[n]},\mathbb{Q})\otimes_{\mathbb{Q}}F$
and the action of the infinite-dimensional Clifford algebra on $F$ is given by geometric correspondences.
1.3. A categorical Boson-Fermion correspondence
The fact that the Bosonic Fock space $B$ is the stable limit of the Fermionic Fock space $F$ is a special case of a categorical Boson-Fermion correspondence which is of independent interest. It reveals an equivalence of certain infinite-dimensional Heisenberg algebra (which we denote as $\mathcal{H}$) representations and graded infinite-dimensional Clifford algebra (which we denote as $\mathcal{C}$) representations. The bridge between $\mathcal{C}$ and $\mathcal{H}$ is an $\mathbb{Q}$-algebra, which we call $\mathcal{E}$:
Definition 1.4 (2.1).
The algebra $\mathcal{E}$ is defined as a $\mathbb{Q}$-algebra generated by $K_{1},K_{-1},E,F$ with relations
$K_{-1}K_{1}=FE=1,\quad K_{-1}E=FK_{1}=0,\quad K_{1}K_{-1}+EF=1.$
The relation between $\mathcal{E}$ with $\mathcal{C}$ is direct: we consider $\{P_{i},Q_{i}\}_{i\in\mathbb{Z}_{>0}}$ such that $P_{1}=EK_{-1},Q_{1}=K_{1}F$ and
$P_{i}=K_{1}P_{i-1}K_{-1}-EP_{i-1}F,\quad Q_{i}=K_{1}Q_{i-1}K_{-1}-EQ_{i-1}F,\quad\text{ if }i>1.$
Then $P_{i},Q_{i}$ satisfy the relations of infinite-dimensional Clifford algebras (see Proposition 2.2 for the details). On the other hand, the Fermionic Fock space $F$ can also be regarded as the highest weight representation of $\mathcal{E}$ generated by a vector $|1\rangle$ and the relations (see Proposition 2.7 for the details)
$F|1\rangle=0,\quad K_{-1}|1\rangle=K_{1}|1\rangle=|1\rangle.$
The relation between the representation theory of $\mathcal{E}$ and $\mathcal{H}$ is more delicate: roughly speaking, given a graded $\mathcal{E}$-representation on
$W=\bigoplus_{i\in\mathbb{Z}}W_{i}$
such that $K_{1}$ has degree $1$, then $K_{1}$ is injective and induces a stable limit
$W_{\infty}=\lim_{n\to\infty}W_{n}$
which is endowed with an $\mathcal{H}$ action if certain admissible property is satisfied. Particularly, we have the following equivalence
Theorem 1.5 (Categorical Boson-Fermion correspondence, Theorem 2.21).
We have a canonical equivalence between the following categories:
  1. (1)
    the abelian category of admissible $\mathcal{H}$-modules;
  2. (2)
    the abelian category of graded admissible $\mathcal{C}$-modules;
  3. (3)
    the abelian category of graded admissible $\mathcal{E}$-modules;
  4. (4)
    the abelian category of $\mathbb{Q}$-vector spaces.
Here the definitions of admissible modules are defined in Definitions 2.8, 2.10 and 2.13.
1.4. Derived Grassmannians and incidence varieties
Due to the categorical Boson-Fermion correspondence, we would like to construct an $\mathcal{E}$-module structure on
$\bigoplus_{m=0}^{\infty}\bigoplus_{n=0}^{\infty}\mathbb{H}^{*}(M^{m}(c_{n}),\mathbb{Q}).$
by incidence varieties. It follows from the incidence variety of the derived Grassmannian which is introduced by Jiang [12]: given a locally free sheaf $\mathcal{I}$ over a variety $X$, the Grassmannian (under the notation of Grothendieck) $\mathcal{I}$, which we denote as $Gr(\mathcal{I},r)$, is the moduli space of quotients
$\mathcal{I}\twoheadrightarrow\mathcal{E}$
where $\mathcal{E}$ is a rank $r$ locally free sheaf on $X$. Jiang [12] generalized the concept of the Grassmannian and allows $\mathcal{I}$ to be a perfect complex, i.e., locally a complex of locally free sheaves. A special case we shall care about is when $\mathcal{I}$ has tor-amplitude $[0,1]$, i.e., it is locally a two-term complex of locally free sheaves. Particularly, in this situation $\mathcal{I}^{\vee}[1]$ is also a tor-amplitude complex.
Fixing integers $d_{-},d_{+}\geq 0$, we consider the derived Grassmannians
$Gr(\mathcal{I},d_{+}),\quad Gr(\mathcal{I}^{\vee}[1],d_{-}).$
In [12], Jiang defined an incidence variety
$Incid_{X}(\mathcal{I},d_{+},d_{-})$
which is the fiber product of $Gr(\mathcal{I},d_{+})$ and $Gr(\mathcal{I}^{\vee}[1],d_{-})$ over $X$ schematically but with a different perfect obstruction theory structure (i.e. derived structure in derived algebraic geometry). The (virtual) fundamental class of $Incid_{X}(\mathcal{I},d_{+},d_{-})$ induces correspondences:
$\phi_{d_{+},d_{-}}:\mathbb{H}^{*}(Gr(\mathcal{I}^{\vee}[1],d_{-}),\mathbb{Q})\to\mathbb{H}^{*}(Gr(\mathcal{I},d_{+}),\mathbb{Q}),$
$\psi_{d_{+},d_{-}}:\mathbb{H}^{*}(Gr(\mathcal{I},d_{+}),\mathbb{Q})\to\mathbb{H}^{*}(Gr(\mathcal{I}^{\vee}[1],d_{-}),\mathbb{Q}).$
Now we come back to the moduli space of perverse coherent sheaves $M^{m}(c_{n})$. Let $\mathcal{I}$ be the universal sheaf over $S^{[n]}\times S$ and
$\mathcal{I}_{o}:=\mathcal{I}|_{S^{[n]}\times o}.$
Then by Nakajima-Yoshioka [22], we have
$Gr_{S^{[n+\frac{l+l^{2}}{2}]}}(\mathcal{I}_{o}^{\vee}[1],l)\cong M^{l}(c_{n})\cong Gr_{S^{[n+\frac{-l+l^{2}}{2}]}}(\mathcal{I}_{o},l).$
Hence the indicence variety of derived Grassmannians induces the maps
$\psi_{l,l-1}:\bigoplus_{n\geq 0}\mathbb{H}^{*}(M^{l}(c_{n}),\mathbb{Q})\to\bigoplus_{n\geq 0}\mathbb{H}^{*}(M^{l-1}(c_{n}),\mathbb{Q}),$
$\phi_{l-1,l}:\bigoplus_{n\geq 0}\mathbb{H}^{*}(M^{l-1}(c_{n}),\mathbb{Q})\to\bigoplus_{n\geq 0}\mathbb{H}^{*}(M^{l}(c_{n}),\mathbb{Q}),$
$\psi_{l,l}:\bigoplus_{n\geq 0}\mathbb{H}^{*}(M^{l}(c_{n}),\mathbb{Q})\to\bigoplus_{n\geq 0}\mathbb{H}^{*}(M^{l}(c_{n-l}),\mathbb{Q}),$
$\phi_{l,l}:\bigoplus_{n\geq 0}\mathbb{H}^{*}(M^{l}(c_{n-l}),\mathbb{Q})\to\bigoplus_{n\geq 0}\mathbb{H}^{*}(M^{l}(c_{n}),\mathbb{Q}).$
The $\mathcal{E}$-representation on the cohomology of moduli space of perverse coherent sheaves follows from the above operators:
Theorem 1.6 (Theorem 6.1).
The operators
$E:=\bigoplus_{l>0}\phi_{l-1,l},\quad F:=\bigoplus_{l>0}\psi_{l-1,l},\quad K_{1}:=\bigoplus_{l\geq 0}\phi_{l,l},\quad K_{-1}:=\bigoplus_{l\geq 0}\psi_{l,l}$
forms an graded admissible $\mathcal{E}$-representation on
$\bigoplus_{m=0}^{\infty}\bigoplus_{n=0}^{\infty}\mathbb{H}^{*}(M^{m}(c_{n}),\mathbb{Q}).$
1.5. Some related work
1.5.1. The affine super-Yangian of $gl(1|1)$
Although not explicitly pointed out in our paper, the representation we constructed is closely related to the quiver Yangian action of Li-Yamazaki [19] and Galakhov-Li-Yamazaki [8]. Here we take an example where $S$ is the equivariant $\mathbb{C}^{2}$ (although it is not projective). Li-Yamazaki [19] defined the affine super-Yangian of $gl(1|1)$, which we denote as
$Y_{\hbar_{1},\hbar_{2}}(\widehat{gl(1|1)}),$
as the $\mathbb{Q}(\hbar_{1},\hbar_{2})$ algebra generated by $e_{n}^{\circ/\bullet}$, $f_{n}^{\circ/\bullet}$, $\psi_{k}^{\circ/\bullet}$, $n\in\mathbb{Z}_{\geq 0}$, $k\in\mathbb{Z}$ with relations:
(1.5) $\begin{array}[]{l}\left\{e^{\circ}_{n},e^{\circ}_{k}\right\}=\left\{f^{\circ}_{n},f^{\circ}_{k}\right\}=\left[\psi^{\circ}_{n},e^{\circ}_{k}\right]=\left[\psi^{\circ}_{n},f^{\circ}_{k}\right]=\left\{e^{\bullet}_{n},e^{\bullet}_{k}\right\}=\left\{f^{\bullet}_{n},f^{\bullet}_{k}\right\}=\left[\psi^{\bullet}_{n},e^{\bullet}_{k}\right]=\left[\psi^{\bullet}_{n},f^{\bullet}_{k}\right]=0\,,\\ \left\{e_{n+2}^{\circ},e_{k}^{\bullet}\right\}-2\left\{e_{n+1}^{\circ},e_{k+1}^{\bullet}\right\}+\left\{e_{n}^{\circ},e_{k+2}^{\bullet}\right\}-\frac{\hbar_{1}^{2}+\hbar_{2}^{2}}{2}\left\{e_{n}^{\circ},e_{k}^{\bullet}\right\}+\frac{\hbar_{1}^{2}-\hbar_{2}^{2}}{2}\left[e_{n}^{\circ},e_{k}^{\bullet}\right]=0\,,\\ \left\{f_{n+2}^{\circ},f_{k}^{\bullet}\right\}-2\left\{f_{n+1}^{\circ},f_{k+1}^{\bullet}\right\}+\left\{f_{n}^{\circ},f_{k+2}^{\bullet}\right\}-\frac{\hbar_{1}^{2}+\hbar_{2}^{2}}{2}\left\{f_{n}^{\circ},f_{k}^{\bullet}\right\}-\frac{\hbar_{1}^{2}-\hbar_{2}^{2}}{2}\left[f_{n}^{\circ},f_{k}^{\bullet}\right]=0\,,\\ \left[\psi_{n+2}^{\circ},e_{k}^{\bullet}\right]-2\left[\psi_{n+1}^{\circ},e_{k+1}^{\bullet}\right]+\left[\psi_{n}^{\circ},e_{k+2}^{\bullet}\right]-\frac{\hbar_{1}^{2}+\hbar_{2}^{2}}{2}\left[\psi_{n}^{\circ},e_{k}^{\bullet}\right]+\frac{\hbar_{1}^{2}-\hbar_{2}^{2}}{2}\left\{\psi_{n}^{\circ},e_{k}^{\bullet}\right\}=0\,,\\ \left[\psi_{n+2}^{\circ},f_{k}^{\bullet}\right]-2\left[\psi_{n+1}^{\circ},f_{k+1}^{\bullet}\right]+\left[\psi_{n}^{\circ},f_{k+2}^{\bullet}\right]-\frac{\hbar_{1}^{2}+\hbar_{2}^{2}}{2}\left[\psi_{n}^{\circ},f_{k}^{\bullet}\right]-\frac{\hbar_{1}^{2}-\hbar_{2}^{2}}{2}\left\{\psi_{n}^{\circ},f_{k}^{\bullet}\right\}=0\,,\\ \left[\psi_{n+2}^{\bullet},e^{\circ}_{k}\right]-2\left[\psi_{n+1}^{\bullet},e_{k+1}^{\circ}\right]+\left[\psi_{n}^{\bullet},e_{k+2}^{\circ}\right]-\frac{\hbar_{1}^{2}+\hbar_{2}^{2}}{2}\left[\psi_{n}^{\bullet},e_{k}^{\circ}\right]-\frac{\hbar_{1}^{2}-\hbar_{2}^{2}}{2}\left\{\psi_{n}^{\bullet},e_{k}^{\circ}\right\}=0\,,\\ \left[\psi_{n+2}^{\bullet},f_{k}^{\circ}\right]-2\left[\psi_{n+1}^{\bullet},f_{k+1}^{\circ}\right]+\left[\psi_{n}^{\bullet},f_{k+2}^{\circ}\right]-\frac{\hbar_{1}^{2}+\hbar_{2}^{2}}{2}\left[\psi_{n}^{\bullet},f_{k}^{\circ}\right]+\frac{\hbar_{1}^{2}-\hbar_{2}^{2}}{2}\left\{\psi_{n}^{\bullet},f_{k}^{\circ}\right\}=0\,,\\ \left\{e^{\circ}_{n},f^{\circ}_{k}\right\}=-\psi^{\circ}_{n+k},\quad\left\{e^{\bullet}_{n},f^{\bullet}_{k}\right\}=-\psi^{\bullet}_{n+k},\quad\left\{e^{\circ}_{n},f^{\bullet}_{k}\right\}=\left\{e^{\bullet}_{n},f^{\circ}_{k}\right\}=0\,,\\ \left[\psi_{n}^{\circ/\bullet},\psi_{k}^{\circ/\bullet}\right]=0\,.\end{array}$
and higher order relations which we will discuss in our future work.
When $S$ is the equivariant $\mathbb{C}^{2}$, the moduli space of perverse coherent sheaves was described as a quiver-like variety by Nakajima-Yoshioka [23]. The affine super-Yangian $Y_{\hbar_{1},\hbar_{2}}(\widehat{gl(1|1)})$ acts on the cohomology of moduli spaces of perverse coherent sheaves by Galakhov-Morozov-Tselousov [9]. The operators in Theorem 1.6 actually come from the action of certain elements in $Y_{\hbar_{1},\hbar_{2}}(\widehat{gl(1|1)})$, which we will discuss in future work.
1.5.2. Cohomological Hall algebra (COHA)
An equivalent definition of the affine super-Yangian $Y_{\hbar_{1},\hbar_{2}}(\widehat{gl(1|1)})$ is the double COHA on the resolved conifold by Soibelmann-Rapcak-Yang-Zhao [26]. Actually, our paper should be regarded as a continuation of Section 7.2.2 of their paper.
By dimension reduction, the affine super-Yangian is also the double COHA of certain one-dimensional objects over a surface, which was defined by Diaconescu-Porta-Sala [6]. A systematic computation of the double COHA will appear very soon [7].
1.5.3. Koseki’s categorification
Another work from which we draw great inspiration is the categorification of the derived category of coherent sheaves on the Hilbert scheme of points by Koseki [16]. To obtain more precise representation-theoretic relations, we apply the semi-orthogonal decomposition theorem of Jiang [14] with explicit Fourier-Mukai kernels.
1.6. Notations in this paper
1.6.1. Derived algebraic geometry
All the varieties and derived schemes in this paper will be over $\mathbb{C}$. For a derived scheme $X$, we will denote $Perf(X)$ as the $\infty$-category of perfect complexes on $X$ and $QCoh(X)$ as the $\infty$-category of quasi-coherent complexes.
1.6.2. Representation theory
[…]
We denote $\mathcal{C}$ as the infinite-dimensional Clifford algebra, i.e., the $\mathbb{Q}$-algebra generated by $\{P_{i},Q_{i}\}_{i\in\mathbb{Z}_{>0}}$ with relations
(1.6) $\{P_{i},P_{j}\}=0,\quad\{Q_{i},Q_{j}\}=0,\quad\{P_{i},Q_{j}\}=\delta_{ij}.$
We also denote $\mathcal{H}$ as the infinite-dimensional Heisenberg algebra, i.e., the $\mathbb{Q}$-algebra generated by $\{H_{i}\}_{i\in\mathbb{Z}-\{0\}}$ with relations
$[H_{i},H_{j}]=i\delta_{i,-j}.$
We denote $B$ as the Bosonic Fermion space, i.e. the polynomial ring of infinitely many variables $\mathbb{Q}[y_{1},y_{2},\cdots]$. It is generated by variables $y_{i},i\in\mathbb{Z}_{>0}$ with the relation
$[y_{i},y_{j}]=0,\forall i,j>0.$
Let $U:=\oplus_{i\in\mathbb{Z}_{>0}}\mathbb{Q}y_{i}$, then $B$ is the symmetric algebra of $U$. We also denote $F$ as the Fermionic Fock space, i.e. the super-polynomial ring of infinitely many variables $\mathbb{Q}[x_{1},x_{2},\cdots]$. It is generated by variables $x_{i},i\in\mathbb{Z}_{>0}$ with the relation
$\{x_{i},x_{j}\}=0,\forall i,j>0.$
Let $V:=\oplus_{i\in\mathbb{Z}_{>0}}\mathbb{Q}x_{i}$, then $F$ is the exterior algebra of $V$.
1.7. Acknowledgements
I want to thank Hiraku Nakajima, Andrei Neguţ, Francesco Sala, Qingyuan Jiang, Jiakang Bao, Dmitry Galakhov, and Henry Liu for many interesting discussions on the subject.
2. A categorical Boson-Fermion correspondence
In this section, we demonstrate an equivalence between specific representations of $\mathcal{C}$ and $\mathcal{H}$ which can be regarded as a categorical version of a Boson-Fermion correspondence. The equivalence is established through a new algebra, denoted as $\mathcal{E}$, generated by $K_{1},K_{-1},E,F$ with the relations
(2.1) $K_{-1}K_{1}=FE=1,\quad K_{-1}E=FK_{1}=0,\quad K_{1}K_{-1}+EF=1.$
It naturally emerges from the orthogonal decomposition of a vector space
$V\cong V_{0}\oplus V_{1}.$
Then the respective operators of projections and embedding
$F:V\to V_{1},\quad E:V_{1}\to V,\quad K_{-1}:V\to V_{0},\quad K_{1}:V_{0}\to V$
satisfy the relations 2.1.
2.1. Basic properties of $\mathcal{E}$ and $\mathcal{E}$-representations
Let
$E_{i}=K_{1}^{i}E,\quad F_{i}=FK_{-1}^{i},\quad i\geq 0.$
Through straightforward calculations, we find that
Lemma 2.1.
The operators $E_{i},F_{i}$ satisfy
(2.2) $\sum_{i=0}^{l}E_{i}F_{i}=1-K_{1}^{l+1}K_{-1}^{l+1},\quad E_{i}F_{i}=K_{1}^{i}K_{-1}^{i}-K_{1}^{i+1}K_{-1}^{i+1};$
and
$K_{1}E_{i}=E_{i+1},\quad K_{-1}E_{i+1}=E_{i},\quad$ $F_{i}K_{-1}=F_{i+1},\quad F_{i+1}K_{1}=F_{i},$
$F_{i}E_{j}=\delta_{i,j},\quad i,j\geq 0$
Let $P_{1}=EK_{-1},Q_{1}=K_{1}F$ and
$P_{i}=K_{1}P_{i-1}K_{-1}-EP_{i-1}F,\quad Q_{i}=K_{1}Q_{i-1}K_{-1}-EQ_{i-1}F,\quad\text{ if }i>1.$
in Section A.1 we prove that
Proposition 2.2.
The elements $\{P_{i},Q_{i}\}_{i\in\mathbb{Z}_{\geq 0}}$ satisfy the relations of $\mathcal{C}$, i.e. 1.6.
Lemma 2.3.
Given an $\mathcal{E}$-module $W$, the actions of $K_{1}$ and $E$ on $W$ are injective, and the actions of $K_{-1}$ and $F$ on $W$ are surjective. Moreover, as vector spaces,
$K_{1}W=\ker(F),\quad EW=\ker(K_{-1}),\quad W=K_{1}W\oplus EW.$
Proof.
The respective injective and surjective properties follow from
(2.3) $K_{-1}K_{1}=FE=1.$
As $K_{-1}E=FK_{1}=0$, we have
$K_{1}W\subset\ker(F),\quad EW\subset\ker(K_{-1}).$
As $K_{1}K_{-1}+EF=1$, we have
$\ker(F)=(K_{1}K_{-1}+EF)\ker(F)=K_{1}K_{-1}\ker(F)\subset K_{1}W.$
Hence $\ker(F)=K_{1}W$ and similarly $EW=\ker(K_{-1})$. Still by the fact that $K_{1}K_{-1}+EF=1$, we have
$W=(K_{1}K_{-1}+EF)W=K_{1}K_{-1}W+EFW\subset K_{1}W+EW,$
$\ker(F)\cap\ker(K_{-1})\subset\ker(K_{1}K_{-1}+EF)=0$
and thus $W=K_{1}W\oplus EW$. ∎
Remark 2.4.
Any non-zero $\mathcal{E}$-module is infinite-dimensional: as $K_{1}$ and $W$ are injective, we have
$W\cong K_{1}W\oplus EW\cong W\oplus W.$
2.2. Highest representation theory of $\mathcal{E}$ and Fermionic Fock space
Definition 2.5.
Given $\lambda\in\mathbb{Q}^{*}$, an $\mathcal{E}$-module $W$ is defined to be the highest weight representation of weight $\lambda$ if there exists $w\in W$ which generates $W$ and
$Fw=0,\quad K_{1}w=\lambda w.$
The Verma module of the highest weight representation of weight $\lambda$ is denoted as $F_{\lambda,\mathcal{E}}$, and the highest weight vector is denoted as 00$\rangle$0 ⟩. We denote $F_{\mathcal{E}}:=F_{1,\mathcal{E}}$.
Given $\lambda\in\mathbb{Q}^{*}$, there is an automorphism of $\mathcal{E}$ which maps
$E\to\lambda E,\quad K_{1}\to\lambda K_{1},\quad F\to\lambda^{-1}F,\quad K_{-1}\to\lambda^{-1}K_{-1}.$
Under this auto equivalence, the representation $F_{\lambda,\mathcal{E}}$ is mapping to $F_{\mathcal{E}}$.
The Fermionic Fock space $F$ has the basis
$x_{i_{1}}\wedge\cdots\wedge x_{i_{k}}$
for any sequence $0<i_{1}<i_{2}<\cdots<i_{k},k\geq 0$. Through straightforward calculations, we find that
Lemma 2.6.
There is a $\mathcal{E}$-representation on the Fermionic Fock space $F$ such that
$K_{1}(x_{j_{1}}\wedge\cdots\wedge x_{j_{k}})=x_{j_{1}+1}\wedge\cdots\wedge x_{j_{k}+1},$
$K_{-1}(x_{j_{1}}\wedge\cdots\wedge x_{j_{k}})=x_{j_{1}-1}\wedge\cdots\wedge x_{j_{k}-1},$
$E(x_{j_{1}}\wedge\cdots\wedge x_{j_{k}})=x_{1}\wedge x_{j_{1}+1}\wedge\cdots\wedge x_{j_{k}+1},$
$F(x_{j_{1}}\wedge\cdots\wedge x_{j_{k}})=\frac{\partial}{\partial x_{1}}(x_{j_{1}-1}\wedge\cdots\wedge x_{j_{k}-1}).$
where we set $x_{0}:=0$. For the induced $\mathcal{C}$-representation on $F$, the action of $P_{i}$ is wedging $x_{i}$ and the action of $Q_{i}$ is $\frac{\partial}{\partial x_{i}}$.
Proposition 2.7.
We have an isomorphism of $\mathcal{E}$-representation
$F_{\mathcal{E}}\cong F$
Proof.
Let $|1\rangle$ be the highest weight vector of $F_{\mathcal{E}}$. Then the map
$|1\rangle\to 1$
induces a morphism from $F_{\mathcal{E}}$ to $F$. We notice that all the vectors
$E_{i_{1}}E_{i_{2}-i_{1}}\cdots E_{i_{n}-i_{n-1}}|1\rangle$
where $0\leq i_{1}\leq i_{2}\leq\cdots\leq i_{n}$ forms a basis of $F_{\mathcal{E}}$, and its image is
$x_{i_{1}}\wedge x_{i_{2}+1}\wedge\cdots\wedge x_{i_{n}+n}.$
Hence the map from $F_{\mathcal{E}}$ to $F$ is an isomorphism.
Now we prove that the Fermionic Fock space $F$ is irreducible as a $\mathcal{E}$-module. By associating all $x_{i}$ with degree $1$, we give a $\mathbb{Z}_{\geq 0}$-grading on $F$. Let $a=\oplus_{i=1}^{N}a_{i}\in F$ such that $a_{i}$ has degree $i$, we take a monomial
$x_{j_{1}}\wedge x_{j_{2}}\wedge\cdots x_{j_{N}}$
in $a_{N}$ and assume its coefficient is $1$. Then
$F_{j_{N}-j_{N-1}-1}\cdots F_{j_{1}-1}a=1.$
Thus the module $\mathcal{E}$ is irreducible. ∎
2.3. Admissible representations and Boson-Fermion correspondence
We consider the category of graded $\mathcal{E}$-modules (resp. $\mathcal{C}$-modules),
$W\cong\bigoplus_{i\in\mathbb{Z}}W_{i}$
such that the degree of $E,F,K_{\pm 1}$ are $1,-1,0$ respectively (resp. the degree of $P_{i}$ is $1$ and $Q_{i}$ is $-1$).
Definition 2.8.
A $\mathbb{Z}_{\geq 0}$-graded $\mathcal{E}$-module
$W\cong\bigoplus_{i=0}^{\infty}W_{i}$
is called admissible if the action of $K_{-1}$ on $W_{i}$ is identity when $i=0$ and locally nilpotent when $i>0$.
Remark 2.9.
By definition we also have $K_{1}|_{W_{0}}=id$.
Definition 2.10.
A graded $\mathcal{C}$-module
$W:=\bigoplus_{i\geq 0}W_{i}$
is defined to be admissible if
  1. (1)
    for any $x\in W$, $Q_{m}x=0$ when $m\gg 0$.
  2. (2)
    We have
    $W_{0}\cong\bigcap_{i>0}ker(Q_{i}).$
Example 2.11.
On the Fermionic Fock space, we associate $x_{i}$ with degree $1$. Then $F$ is both a graded admissible $\mathcal{E}$-module and a graded admissible $\mathcal{C}$-module.
Example 2.12.
We consider the Bosonic Fock space
$B:=\bigoplus_{i=0}^{\infty}V\cong\mathbb{Q}[y_{1},y_{2},\cdots].$
It has the $\mathcal{H}$-representation structure by
$\begin{cases}H_{k}\to y_{k},\quad k>0,\\ H_{k}\to\frac{\partial}{\partial y_{-k}},\quad k<0.\end{cases}$
For a sequence of non-negative integers $\lambda=(n_{1},n_{2},\cdots)$ such that $n_{m}=0$ when $m\gg 0$, we denote $y_{\lambda}:=\prod_{i=1}^{\infty}y_{i}^{n_{i}}$, which is a monomial of finite variables.
We also introduce the admissible $\mathcal{H}$-representations:
Definition 2.13.
Given a sequence of non-negative integers $\lambda=(n_{1},n_{2},\cdots)$ such that $n_{m}=0$ when $m\gg 0$, we denote
$|\lambda|=\sum_{i=1}^{\infty}in_{i},\quad H_{\lambda}=\prod_{i=1}^{\infty}H_{i}^{n_{i}},\quad H_{-\lambda}=\prod_{i=1}^{\infty}H_{-i}^{n_{i}}.$
An $\mathcal{H}$-representation $W$ is called admissible, if for any $x\in W$ there exists an integer $N>0$ such that $H_{-\lambda}x=0$ for any $|\lambda|>N$.
Remark 2.14.
The $\mathcal{H}$-representation $B$ is admissible.
In Sections 2.4, 2.5 and 2.6, we establish the following equivalence
Proposition 2.15.
There is an equivalence among the abelian category of admissible $\mathcal{H}$-modules, graded admissible $\mathcal{C}$-modules, and graded admissible $\mathcal{E}$-modules.
2.4. Graded $\mathcal{E}$-modules
We first establish the equivalence between the category admissible graded $\mathcal{E}$-modules and graded $\mathcal{C}$-modules. For any $\mathcal{E}$-module $W$, the morphsim from $\mathcal{C}$ to $\mathcal{E}$ induces a $\mathcal{C}$-representation on $W$.
Proposition 2.16.
If $W$ is an admissible graded $\mathcal{E}$-module, the induced $\mathcal{C}$-module on $W$ is also admissible.
Proof.
First we prove that for any $x$ in $W$, $Q_{m}x=0$ when $m\gg 0$ by induction. If $x\in W_{0}$, it follows from the fact that $Q_{1}x=0$ and $Fx=0$. Assuming it is true for $x\in W_{i}$, then for $x\in W_{i+1}$, we have that for some $N>0$
$x=\sum_{i=1}^{N}E_{i}F_{i}x$
By A.5 and induction, we have
$Q_{m}E_{i}F_{i}x=E_{i}Q_{m-i}F_{i}x=0$
when $m\gg 0$. Hence $Q_{m}x=0$ for $m\gg 0$.
Now we prove that
$W_{0}\cong\bigcap_{i>0}ker(Q_{i}).$
First we notice that for any non-zero $x\in\oplus_{i>0}W_{i}$, there exists $M\geq 0$ such $F_{M}x\neq 0$. Otherwise, by 2.2, for any $M>0$
$K_{-1}^{M}x=x$
which contradicts the assumption of admissibility. Let $M_{0}$ be the smallest integer such that $F_{M_{0}}x\neq 0$. Then by induction, we have
$Q_{M_{0}+1}x=K_{1}Q_{M_{0}}K_{-1}x=\cdots=K_{1}^{M_{0}}Q_{1}K_{-1}^{M_{0}}x=K_{1}^{M_{0}+1}F_{M_{0}}x.$
which is non-zero as $K_{1}$ is injective. ∎
On the other hand, it is more complicated to construct a $\mathcal{E}$-representation from a graded $\mathcal{C}$-module, which we will leave to Proposition A.4.
2.5. Filtration and Fermionization
In this subsection we give the functor from admissible $\mathcal{H}$-representations to graded admissible $\mathcal{E}$-representations: let $W_{\infty}$ be an admissible $\mathcal{H}$-representation, we consider
$W_{i}:=\bigcap_{l>i}ker(H_{l}).$
It satisfies that $W_{i}\subset W_{j}$ for $i\leq j$ and
$W_{\infty}=\bigcup_{i\geq 0}W_{i}.$
Hence it forms a filtration of $W$. For any $k>0$, we consider operators
$\mathcal{L}_{k}:=\sum_{i=1}^{\infty}\frac{-1}{k^{i}i!}H_{k}^{i-1}H_{-k}^{i}$
The admissible property of $W_{\infty}$ makes sure that the action of $\mathcal{L}_{k}$ on $W_{\infty}$ is well defined. By the commutative property of the Heisenberg algebra, we have
(2.4) $H_{k}W_{k}\subset W_{k},\quad\mathcal{L}_{k}W_{k}\subset W_{k}.$
In Section A.3, we will prove that
Proposition 2.17.
We have
(2.5) $(Id-H_{k}\mathcal{L}_{k})W_{k}\subset W_{k-1}.$
Let $E:W_{k}\to W_{k+1}$ be the embedding operator, and $F:W_{k+1}\to W_{k}$ be the operator induced by $Id-H_{k}\mathcal{L}_{k}$. Moreover, we consider the operators
$K_{1}:=\bigoplus_{k\geq 0}H_{k},\quad K_{-1}:=\bigoplus_{k\geq 0}\mathcal{L}_{k},$
where $H_{0}=\mathcal{L}_{0}:=id$. Then $W$ is an graded admissible $\mathcal{E}$-module under the action of $E,F,K_{1},K_{-1}$.
2.6. Stable limit and Bosonization
The functor from graded admissible $\mathcal{E}$-modules to admissible $\mathcal{H}$-modules is constructed by taking the stable limit:
Definition 2.18.
Given a graded admissible $\mathcal{E}$-module
$W:=\bigoplus_{i\geq 0}W_{i}$
we consider the stable limit,
$W_{\infty}:=\lim_{n\to\infty}W_{i}$
where the morphism form $W_{i}$ to $W_{i+1}$ is $K_{1}$.
By induction, there exists a unique series of operators $\{H_{i,j}\}|_{-i\leq j\leq i,j\neq 0}$ which acts on $W_{i}$ such that
$H_{i,i}:=K_{1}|_{W_{i}},\quad H_{-i,i}:=-i\sum_{l=0}^{\infty}(K_{1})^{l}K_{-1}^{l+1}|_{W_{i}}.$
and if $|j|\neq i$
$H_{i,j}:=\sum_{l=0}^{\infty}E_{l}H_{i-1,j}F_{l}.$
Moreover, as
$H_{i+1,j}E=\sum_{l=0}^{\infty}E_{l}H_{i,j}F_{l}E=\sum_{l=0}^{\infty}E_{l}H_{i,j}\delta_{0,l}=E_{l}H_{i,j},$
for any $l\neq 0$, the operators $\{H_{i,l}\}_{i\geq 0}$ converges to an operator $H_{l}$ on $W_{\infty}$. In the Section A.4, we will prove that
Proposition 2.19.
The operators $H_{l}|_{l\in\mathbb{Z}-\{0\}}$ forms an admissible $\mathcal{H}$-representation on $W_{\infty}$.
Proof of Proposition 2.15.
It follows from Propositions 2.16, A.4, 2.17 and 2.19. ∎
2.7. The Boson-Fermion correspondence
The Fermion Fock space $F$ is a graded admissible $\mathcal{E}$-module and the Bosonic Fock space $B$ is an admissible $\mathbb{H}$-module. We have that
Proposition 2.20.
The stable limit of the Fermionic Fock space $F$ is the Bosonic Fock space $B$.
Proof.
We construct the mapping from $\rho:F\to B$
$\rho(x_{i_{1}}\wedge x_{i_{2}}\cdots\wedge x_{i_{n}}):=y_{1}^{i_{n}-i_{n-1}-1}y_{2}^{i_{n-1}-i_{n-2}-1}\cdots y_{n}^{i_{1}-1}.$
It is easy to see the following properties of $\rho$
  1. (1)
    $\rho$ is invariant under $K_{1}$, i.e. for any $x\in F$,
    $\rho(K_{1}x)=\rho(x).$
  2. (2)
    For any $i\geq 0$, $\rho$ induces an isomorphism of vector spaces between $F_{i}$ and
    $B_{i}\cong\mathbb{Q}[y_{1},\cdots,y_{i}].$
hence by taking the limit $i\to\infty$, we prove Proposition 2.20. ∎
Theorem 2.21 (Categorical Boson-Fermion Correspondence).
By tensoring the Fermionic or Bosonic Fock space, it induces an equivalence from the abelian category of vector spaces to the abelian category of graded admissible $\mathcal{C}$-modules or admissible $\mathcal{H}$-modules.
By Proposition 2.15, Theorem 2.21 follows from the following proposition:
Proposition 2.22.
For any graded admissible $\mathcal{C}$-module $W$,
$W\cong F\otimes_{\mathbb{Q}}W_{0}.$
Proof.
For any $k>0$, we consider
$\bar{W}^{k}:=\bigcap_{l>k}ker(Q_{l})$
which is a $\mathbb{Z}_{\geq 0}$-grade vector space and
$W=\lim_{k\to\infty}\bar{W}^{k}.$
We notice that
$F\cong\lim_{k\to\infty}\wedge_{i=1}^{k}(\mathbb{Q}\oplus\mathbb{Q}x_{i}).$
Hence to prove Proposition 2.22, we only need to prove that for any $k\geq 0$,
$\bar{W}^{k}\cong W_{0}\otimes_{\mathbb{Q}}\bigotimes_{i=1}^{k}\wedge^{i}(\mathbb{Q}\oplus\mathbb{Q}x_{i}).$
Moreover, by induction, we only need to prove that
$\bar{W}^{k+1}\cong\bar{W}^{k}\otimes_{\mathbb{Q}}(\mathbb{Q}\oplus\mathbb{Q}x_{k+1}).$
as $\bar{W}^{0}\cong W_{0}$. It follows from the action of the Clifford algebra generated by $P_{k+1}$ and $Q_{k+1}$. ∎
3. Derived categories of derived Grassmannians
Given a variety $X$, a locally free sheaf $\mathcal{V}$ and an integer $r\geq 0$, the Grassmannian (on Grothendieck’s notation) $Gr_{X}(\mathcal{V},r)$ is defined as the moduli space of quotients
$\mathcal{V}\twoheadrightarrow\mathcal{E}$
where $\mathcal{E}$ is a locally free sheaf of rank $r$.
Jiang [13, 14, 12] extended the definition of Grassmannians and flag varieties such that $\mathcal{V}$ can be any perfect complex. Particularly, when $\mathcal{V}$ is a tor-amplitude $[0,1]$ perfect complex, i.e. locally a two-term complexes of locally free sheaves
$\mathcal{V}\cong\{W\to V\}$
the derived Grassmannian is quasi-smooth (i.e. the locally complete intersection scheme in the derived algebraic geometry setting). Moreover, Jiang [14] strengthens the result of Toda [28] for the semi-orthogonal decomposition of derived Grassmannians by given Fourier-Mukai kernels. In this section, we review several main results of Jiang [13, 14, 12].
In this section, we will always assume that $X$ is a derived scheme and $\mathcal{E}$ is a Tor-amplitude $[0,1]$ perfect complex on $X$ of rank $e$.
3.1. Derived Grassmannians of a Tor-amplitude $[0,1]$ complex
Definition 3.1 (Definition 4.3 of [12]).
Given an integer $d\geq 0$, we define the derived Grassmannian
$Gr_{X}(\mathcal{E},d)$
as the derived stack which parametrized the moduli space of quotients
$\{\mathcal{E}\twoheadrightarrow\mathcal{V}\}$
where $\mathcal{V}$ is a rank $d$ locally free sheaf. We define
$pr_{\mathcal{E},d}:Gr_{X}(\mathcal{E},d)\to X$
as the projection morphism.
The concept of the quasi-smooth morphisms (schemes) is a generalization of the locally complete intersection morphisms (schemes) in derived algebraic geometry:
Definition 3.2.
A morphism of derived schemes $f:X\to Y$ is defined to be quasi-smooth if the cotangent complex $L_{f}:=L_{X/Y}$ is of tor-amplitude $[0,1]$. A derived scheme $X$ is defined to be quasi-smooth if the morphism from $X$ to $\mathrm{Spec}(\mathbb{C})$ is quasi-smooth. For a quasi-smooth morphism $f$ (resp. scheme $X$), we define its virtual dimension, denoted as
$vdim(f),\quad(\text{resp. }vdim(X))$
as the rank of the cotangent complex $L_{f}$ (resp. $L_{X}$). The relative canonical bundle $K_{f}$ (resp. $K_{X}$) is defined as the determinant of $L_{f}$ (resp. $L_{X}$). (We refer to [27] or [13] for the determinant of a tor-amplitude $[0,1]$ complex).
Remark 3.3.
Given a quasi-smooth derived scheme $X$ over $\mathbb{C}$, we say that $X$ is classical if it has a trivial derived structure, i.e. $\pi_{0}(X)\cong X$ where $\pi_{0}(X)$ is the underlying scheme. By Corollary 2.11 of [1] (which we also discussed in the Appendix A of [30]), $X$ is classical if and only if
$vdim(X)=dim(\pi_{0}(X)).$
and $X$ is always a locally complete intersection variety if $X$ is classical.
Theorem 3.4 ([12]).
The derived Grassmannian $Gr_{X}(\mathcal{E},d)$ is a proper derived scheme over $X$ and the projection map
$pr_{\mathcal{E},d}:Gr_{X}(\mathcal{E},d)\to X$
is quasi-smooth with relative virtual dimension is $d(rank(\mathcal{E})-d)$.
For a proper quasi-smooth morphism $f:X\to Y$, the derived pullback and pushforward are all well-defined in the $\infty$-category of perfect complexes
$f^{*}:Perf(Y)\to Perf(X),\quad f_{*}:Perf(X)\to Perf(Y).$
Let $f_{!}$ be the left-adjoint functor $f^{*}$, and $f^{!}$ by the right-adjoint functor of $f_{*}$. By the Grothendieck duality, we have
(3.1) $f_{!}\cong f_{*}(-\otimes K_{Y/X})[vdim(f)],\quad f^{!}\cong(-\otimes K_{Y/X})\circ f^{*}[vdim(f)].$
3.2. The incidence varieties
We notice that if $\mathcal{E}$ has tor-amplitude $[0,1]$, the shifted dual $\mathcal{E}^{\vee}[1]$ also has tor amplitude $[0,1]$.
Definition 3.5 (Incidence correspondence, Definition 2.7 of [14]).
Given $(d_{+},d_{-})\in\mathbb{Z}_{\geq 0}^{2}$, the incidence scheme $Inc_{X}(\mathcal{E},d_{+},d_{-})$ can be defined in two ways, which are equivalent due to [12]:
  1. (1)
    Over the projection morphism $pr_{+}:Gr_{X}(\mathcal{E},d_{+})\to X$, let $\mathcal{E}_{+}$ be the fiber of the universal quotient
    $\theta_{+}:pr^{*}_{+}\mathcal{E}\to\mathcal{V}^{+}.$
    we define
    $Inc_{X}(\mathcal{E},d_{+},d_{-}):=Gr_{Gr_{X}(\mathcal{E},d+)}(\mathcal{E}_{+}^{\vee}[1],d_{-}).$
  2. (2)
    Over the projection morphism $pr_{-}:Gr_{X}(\mathcal{E}^{\vee}[1],d_{-})\to X$, let $\mathcal{E}_{-}$ be the fiber of the universal quotient
    $\theta_{-}:pr^{*}_{-}\mathcal{E}^{\vee}[1]\to\mathcal{V}^{-\vee}.$
    we define
    $Inc_{X}(\mathcal{E},d_{+},d_{-}):=Gr_{Gr_{X}(\mathcal{E}^{\vee}[1],d_{-})}(\mathcal{E}_{-}^{\vee}[1],d_{+}).$
The above definitions induce canonical projection morphisms:
$r^{+}_{d_{+},d_{-}}:Inc_{X}(\mathcal{E},d_{+},d_{-})\to Gr_{X}(\mathcal{E},d_{+}),\quad r^{-}_{d_{+},d_{-}}:Inc_{X}(\mathcal{E},d_{+},d_{-})\to Gr_{X}(\mathcal{E}^{\vee}[-1],d_{-}).$
which we will also abberivate as $r^{+},r^{-}$ respectively.
By Theorem 3.4, $r^{+}$ and $r^{-}$ are both quasi-smooth and proper, and the relative virtual dimensions are
$vdim(r^{+})=d_{-}(-e+d_{+}-d_{-}),\quad vdim(r^{-})=d_{+}(e+d_{-}-d_{+}).$
The virtual dimension of the incidence variety $Inc_{X}(\mathcal{E},d_{+},d_{-})$ over $X$ is
$d_{-}(-e-d_{-})+d_{+}(e-d_{+})+d_{-}d_{+}.$
Hence the functor
$r^{+}_{*}r^{-*}:Perf(Gr_{X}(\mathcal{E}^{\vee}[-1],d_{-}))\to Perf(Gr_{X}(\mathcal{E},d_{+}))$
is represented by the Fourier-Mukai transform
$\Phi_{d_{+},d_{-}}:=(r^{+}\times r^{-})_{*}(\mathcal{O}_{Inc_{X,d_{+},d_{-}}})\in QCoh(Gr_{X}(\mathcal{E}^{\vee}[-1],d_{-})\times Gr_{X}(\mathcal{E},d_{+})),$
Moreover, let $\Psi_{d_{+},d_{-}}^{L}$ and $\Psi_{d_{+},d_{-}}$ be the left and adjoint functor of $\Phi_{d_{+},d_{-}}$. By the Grothendieck duality 3.1, we have
$\Psi_{d_{+},d_{-}}^{L}\cong(r^{-}\times r^{+})_{*}(K_{r^{-}}[vdim(r^{-})]),\quad\Psi_{d_{+},d_{-}}\cong(r^{-}\times r^{+})_{*}(K_{r^{+}}[vdim(r^{+})])$
as the Fourier-Mukai kernels.
Remark 3.6.
We will not calculate $K_{r^{-}}$ and $K_{r^{+}}$ explicitly in this paper as we do not need it. The reader can refer to [13] for the explicit formula. However, we should point out that they are generated by tautological complexes over the incidence variety.
3.3. Semi-orthogonal decomposition and decomposition of the diagonal
When $e:=rank(\mathcal{E})\geq 0$ is positive, by Toda [28] and Jiang [14], for any $d\geq 0$ the category
$Perf(Gr_{X}(\mathcal{E},d))$
has a semi-orthogonal decomposition by copies of
$Perf(Gr_{X}(\mathcal{E}^{\vee}[1],d^{\prime})),\quad max\{d-e,0\}\leq d^{\prime}\leq d.$
Moreover, Jiang [14] proved that the functors are given as Fourier-Mukai transforms through the incidence varieties. Now we introduce a special case when $e=1$.
Theorem 3.7 ([14]).
If $e=1$, then for any $d>0$, the functors $\Phi_{d,d}$ and $\Phi_{d.d-1}$ are all fully faithful and we have the semi-orthogonal decomposition
$Perf(Gr_{X}(\mathcal{E},d))=<Im(\Phi_{d,d-1}),Im(\Phi_{d,d})>.$
Corollary 3.8.
Let $\Psi_{d,d-1}$ be the right adjoint functor of $\Phi_{d,d-1}$ and $\Psi_{d,d}^{L}$ be the left-adjoint fucntor of $\Phi_{d,d}$. Then there exists a canonical triangle of Fourier-Mukai transforms
$\Phi_{d,d-1}\Psi_{d,d-1}\to\mathcal{O}_{\Delta}\to\Phi_{d,d}\Psi_{d,d}^{L}$
where $\Delta$ is the diagonal embedding of $Gr_{X}(\mathcal{E},d)$. Moreover, we have
$\Psi_{d,d-1}\Phi_{d,d-1}\cong\mathcal{O}_{\Delta},\quad\Psi_{d,d}^{L}\Phi_{d,d}\cong\mathcal{O}_{\Delta},\quad\Psi_{d,d-1}\Phi_{d,d}\cong 0,\quad\Psi_{d,d}^{L}\Phi_{d,d-1}\cong 0.$
Proof.
trans We only need to show natural transformations of functors are induced by morphisms of Fourier-Mukai kernels. As all the settings can be naturally lifted to $\infty$-categories, it follows from Theorem 1.2(2) of [3]. ∎
From now on, we denote
$\Psi_{d,d-1}:=\Psi_{d,d-1}^{R},\quad\Psi_{d,d}:=\Psi_{d,d}^{L}.$
Corollary 3.9.
Let $[\Psi_{d,d-1}]$, $[\Psi_{d,d}]$, $[\Psi_{d,d-1}]$, $[\Psi_{d,d}]$ be the corresponding correspondences at the level of Grothendieck group of perfect complexes. Then we have
(3.2) $[\Psi_{d,d-1}][\Phi_{d,d-1}]=id=[\Psi_{d,d}][\Phi_{d,d}]\quad[\Psi_{d,d-1}][\Phi_{d,d}]=0=[\Psi_{d,d}][\Phi_{d,d-1}],$
and
(3.3) $[\Phi_{d,d-1}][\Psi_{d,d-1}]+[\Phi_{d,d}][\Psi_{d,d}]=id.$
4. Mukai pairing, K-theoretic and cohomological correspondences
Let $\mathbb{H}$ be a cohomology theory other than the Grothendieck groups of coherent sheaves. In this section, we always assume that $X$ and $Y$ are smooth projective varieties. Given an element in the Grothendieck group of coherent sheaves $f\in K(X\times Y)$, one can define the K-theoretic correspondence
$\Gamma_{X\to Y}^{\mu}:K(X)\to K(Y),\quad\Gamma_{X\to Y}^{\mu}(\cdot)=\pi_{Y,*}(\pi_{X}^{*}(\cdot)\otimes\mu).$
Similarly, for any element $u\in\mathbb{H}^{*}(X\times Y,\mathbb{Q})$, one can define the cohomological correspondence
$\gamma_{X\to Y}^{u}:\mathbb{H}^{*}(X,\mathbb{Q})\to\mathbb{H}^{*}(Y,\mathbb{Q}),\quad\gamma_{X\to Y}^{u}(\cdot)=\pi_{Y*}(\pi_{X}^{*}(\cdot).\mu).$
Here $\mathbb{H}^{*}$ can be one of the Chow groups $CH^{*}$, the singular cohomology $H^{*}$, the Hodge cohomology $H^{*,*}$ or the Hochschild homology $HH_{*}$.
One may expect that two transforms can be naturally compared by the Chern character, i.e. we have a commutative diagram
(4.1) ${K(X)}$${K(Y)}$${\mathbb{H}^{*}(X,\mathbb{Q})}$${\mathbb{H}^{*}(Y,\mathbb{Q})}$$\scriptstyle{\Gamma_{X\to Y}^{\mu}}$$\scriptstyle{ch}$$\scriptstyle{ch}$$\scriptstyle{\gamma_{X\to Y}^{ch(\mu)}}$
where $ch$ is the Chern character. Unfortunately, 4.1 in general is not true due to the Grothendieck-Riemann-Roch theorem. Caldararu [4] found the correct way to modify the above diagram by generalizing the Mukai vectors, which we will recall in this section.
4.1. K-theoretic and cohomological correspondences
By the cohomological grading, we have a decomposition based on the odd or even degree
$H^{*}(X,\mathbb{Q})\cong H^{even}(X,\mathbb{Q})\oplus H^{odd}(X,\mathbb{Q})$
where $H^{even}(X,\mathbb{Q})$ is a commutative ring. By Section 2.4 of [4], there exists a unique formal series expansion such that
$\sqrt{1}=1,\quad\sqrt{uv}=\sqrt{u}\sqrt{v},\quad u=(\sqrt{u})^{2}$
for every smooth projective variety $X$ and any $u,v\in H^{even}(X,\mathbb{Q})$ (resp. $CH^{*}(X,\mathbb{Q})$ and $HH_{0}(X)$) with constant term $1$. We define $\hat{A}_{X}\in\mathbb{H}^{*}(X,\mathbb{Q})$ as
$\hat{A}_{X}:=td_{X}\cdot\sqrt{ch(K_{X})}$
where $K_{X}$ is the canonical line bundle of $X$ and $td_{X}$ is the Todd class of $X$.
Definition 4.1 (Definition 2.1 of [4]).
The directed Mukai vector of an element $\mu\in K(X\times Y)$ is defined by
$v(\mu,X\to Y):=ch(\mu)\cdot\sqrt{td_{X\times Y}}\cdot\sqrt[4]{ch(K_{Y})/ch(K_{X})}\in\mathbb{H}^{*}(X\times Y)$
By taking the first space to be a point we obtain the definition of the Mukai vector of an object $\alpha\in K(X)$ as
$v(\alpha)=ch(\alpha)\sqrt{\hat{A}_{X}}\in\mathbb{H}^{*}(X).$
Definition 4.2.
We define a map $(-)_{*}$ from the $K$-theoretic correspondences on $K(X\times Y)$ to cohomological correspondences $\mathbb{H}^{*}(X\times Y,\mathbb{Q})$ by
$\Gamma_{X\to Y*}^{\mu}:=\gamma_{X\to Y}^{v(\mu)}.$
Proposition 4.3 (Section 2.3 of [4]).
The map $(-)_{*}$ is linear with the following properties:
  1. (1)
    Given correspondences $\mu:K(X)\to K(Y)$, $\nu:K(Y)\to K(Z)$, we have
    $\nu_{*}\circ\mu_{*}\cong\nu_{*}\circ\mu_{*};$
  2. (2)
    $id_{*}\cong id$;
  3. (3)
    The following commutative diagram commutes:
    (4.2) ${K(X)}$${K(Y)}$${\mathbb{H}^{*}(X,\mathbb{Q})}$${\mathbb{H}^{*}(Y,\mathbb{Q})}$$\scriptstyle{\Gamma_{X\to Y}^{\mu}}$$\scriptstyle{v}$$\scriptstyle{v}$$\scriptstyle{\gamma_{X\to Y}^{v(\mu)}}$
4.2. Cohomological correspondences of derived Grassmannians
Now we come back to the setting of Section 3.2. Let $X$ be a smooth proper variety with a tor-amplitude $[0,1]$ perfect complex $\mathcal{E}$. We assume that for all $d\geq 0$, the derived Grassmannian
$Gr_{X}(\mathcal{E},d),\quad Gr_{X}(\mathcal{E}^{\vee}[1],d)$
are (classical) smooth varieties. We apply the Caldararu’s construction for $[\Phi_{d_{+},d_{-}}]$ and by 3.2 and 3.3 and have
Corollary 4.4.
The cohomological correspondences $[\Phi_{d,d-1}]_{*},[\Psi_{d,{d-1}}]_{*},[\Phi_{d,d}]_{*},[\Psi_{d,d}]_{*}$ satisfy the relations 2.1, if we denote
$E:=[\Phi_{d,d-1}]_{*},\quad F:=[\Psi_{d,d-1}]_{*},\quad K_{1}:=[\Phi_{d,d}]_{*},\quad K_{-1}:=[\Phi_{d,d}]_{*}.$
The correspondences in Corollary 4.4 are homogeneous with degree 00 for the Hochschild homology, but not for the singular cohomology or Chow groups and we need to modify them. Recall that the incidence variety
$Inc_{X}(\mathcal{E},d_{+},d_{-})$
is quasi-smooth with virtual dimension $d_{0}=dim(X)+(1-d_{+})d_{+}-(d_{-}+1)d_{-}+d_{-}d_{+}$. By Ciocan-Fontanine and Kapranov [5], the derived structure induces a perfect obstruction theory on the classical scheme of $Inc(X,d_{+},d_{-})$ and thus induces a virtual fundamental class in the sense of Li-Tian [18] or Behrend-Fantechi [2]:
$Vir_{d_{+},d_{-}}:=[Inc_{X}(\mathcal{E},d_{+},d_{-})]^{vir}\in CH_{d_{0}}(Inc_{X}(\mathcal{E},d_{+},d_{-})).$
Now we define cohomological correspondences $\psi_{d,d-1},\psi_{d,d},\phi_{d,d-1},\phi_{d,d}$ such that
  1. (1)
    $\phi_{d,d}$ is induced by $Vir_{d,d}$, and $\psi_{d,d}$ is induced by $(-1)^{d}Vir_{d,d}$ but in a different direction.
  2. (2)
    both $\phi_{d,d-1}$ and $\psi_{d,d-1}$ are induced by $Vir_{d,d-1}$ but in two different directions.
Proposition 4.5.
The operators $\phi_{d,d-1}$ and $\psi_{d,d-1}$ preserved the cohomological degree and the operators $\phi_{d,d}$ and $\psi_{d,d}$ shifted the Chow group (resp. cohomological, Hodge) degree by $d$ (resp. $2d$, $(d,d)$) and $-d$ (resp. $-2d$, $(-d,-d)$). Moreover, the cohomological correspondences $\phi_{d,d-1},\psi_{d,s-1},\phi_{d,d},\psi_{d,d}$ also satisfy the relations 2.1 if we denote
$E:=\phi_{d,d-1},\quad F:=\psi_{d,d-1},\quad K_{1}:=\phi_{d,d},\quad K_{-1}:=\phi_{d,d}.$
Proof.
First, we notice that all the constant terms of roots of the Todd class and Chern classes of line bundles are $1$. By the virtual Grothendieck-Riemann-Roch theorem (i.e. Theorem 3.23 and Corollary 3.25 of [15]), we also have
$ch(\mathcal{O}_{Inc_{X}(\mathcal{E},d_{+},d_{-})})=Vir_{d_{+},d_{-}}+C_{d_{+},d_{-}}$
where $C_{d_{+},d_{-}}$ have higher degrees than $Vir_{d_{+},d_{-}}$. Thus we have $[\Phi_{d,d}]_{*}-\phi_{d,d}$ has a higher cohomological degree than $\phi_{d,d}$. Similar computations also holds for $[\Phi_{d,d-1}]_{*}-\phi_{d,d-1},[\Psi_{d,d}]_{*}-\psi_{d,d}$ and $[\Psi_{d,d-1}]_{*}-\psi_{d,d-1}$. Thus by checking the cohomological grading, the relations 2.1 for $\phi_{d,d-1},\psi_{d,d-1},\phi_{d,d},\psi_{d,d}$ follows from Corollary 4.4. ∎
In a companion paper [29], we will prove that
Proposition 4.6.
The incidence variety
$Inc_{X}(\mathcal{E},d_{+},d_{-})$
is classical and
$Vir_{d_{+},d_{-}}:=[Inc_{X}(\mathcal{E},d_{+},d_{-})]\in CH_{d_{0}}(Inc_{X}(\mathcal{E},d_{+},d_{-})).$
5. Perverse Coherent Sheaves and their moduli spaces
Let $S$ be a smooth projective surface. We denote by
$p:\hat{S}\to S$
the blow up of $S$ at a closed point $o\in S$. We denote $C:=p^{-1}(o)$ as the exceptional (rational) curve. We denote
$\mathcal{O}_{C}(m):=\mathcal{O}(-mC)|_{C}$
which is the unique degree $m$ line bundle on the rational curve $C$ for any $m\in\mathbb{Z}$. Given an ample line bundle $H$ on $S$, Nakajima-Yoshioka defined the perverse coherent sheaf as
Definition 5.1.
A stable perverse coherent sheaf $E$ on $\hat{S}$ with respect to $H$ is a coherent sheaf such that
$Hom(E,\mathcal{O}_{C}(-1))=0,\quad Hom(\mathcal{O}_{C},E)=0$
and $p_{*}E$ is Gieseker stable with respect to $H$.
Given a coherent sheaf $E$ on $\hat{S}$ and an integer $m$, we define $E$ to be $m$-stable if $E\otimes\mathcal{O}(-mc)$ is a perverse stable sheaf.
Remark 5.2.
When the rank of $E$ is $1$, $p_{*}E$ is stable with respect to $H$ if and only if $p_{*}E$ is torsion-free, and hence independent with the choice of $H$.
By the work of Nakajima-Yoshioka [23, 22, 24], the stable perverse coherent sheaf builds a bridge between the Hilbert scheme of points on $S$ and $\hat{S}$ (and more generally the moduli space of stable sheaves), which we recall in this section.
5.1. Moduli space of perverse coherent sheaves
We recall the decomposition of cohomology groups:
$H^{*}(\hat{S},\mathbb{Q})=p^{*}(H^{*}(\hat{S},\mathbb{Q}))\oplus\mathbb{Q}[C]$
where $[C]\in H^{1,1}(\hat{S})\cap H^{2}(\hat{S},\mathbb{Q})$ is the fundamental class of $C$ satisfying $[C]^{2}=-1$.
Definition 5.3.
Given Chern characters $\hat{c}\in H^{*}(\hat{S},\mathbb{Q})$, we denote $M^{m}(\hat{c})$ as the moduli space of $m$-stable coherent sheaves on $\hat{S}$ with Chern character $\hat{c}$. Particularly, given integers $l,n$, we denote
$M^{m}(l,n):=M^{m}([\hat{S}]-l[C]+(-n-\frac{l^{2}}{2})[o]),$
i.e. it consists of $m$-stable perverse coherent sheaves $E$ such that
$rank(E)=1,\quad c_{1}(E)=-l[C],\quad c_{2}(E)=n$
By tensoring $\mathcal{O}(-mC)$, we have a canonical isomorphism
$M^{m}(\hat{c})\cong M^{0}(\hat{c}e^{-m[C]}),\quad M^{m}(l,n)\cong M^{0}(l+m,n)$
Lemma 5.4 (Lemma 3.3 of [16] and Proposition 3.37 of [22] ).
For any integers $l,n$, the variety $M^{0}(l,n)$ is either empty or a smooth projective variety of $2n$ dimension. Particularly, fixing $n$, we have
$M^{0}(0,n)\cong S^{[n]},\quad M^{0}(l,n)\cong\hat{S}^{[n]}\text{ if }l\gg 0.$
5.2. Moduli spaces as derived Grassmannians
We first recall the cohomological properties of stable coherent sheaves
Proposition 5.5 (Proposition 3.1 of [17] and Lemma 3.4 of [22]).
For a stable perverse coherent sheaf $E$ on $\hat{S}$, we have
$R^{i}p_{*}E(aC)=0,\quad a=0,1,\quad i\geq 1$
and moreover
$Rp_{*}E(aC)=p_{*}E(aC),\quad a=0,1$
are all stable coherent sheaves with respect to $H$.
For any stable perverse coherent sheaf $E$ such that
$rank(E)=1,\quad c_{1}(E)=-l[C],\quad c_{2}(E)=n,$
by Proposition 5.5
$p_{*}E,\text{ and }p_{*}E(C)$
are all stable and by the Grothendieck-Riemann-Roch theorem we have
$rank(p_{*}E)=1,\quad c_{1}(p_{*}E)=0,\quad c_{2}(p_{*}E)=(n+\frac{l+l^{2}}{2}).$
$rank(p_{*}E(C))=1,\quad c_{1}(p_{*}E(C))=0,\quad c_{2}(p_{*}E(C))=(n+\frac{-l+l^{2}}{2}).$
It induces morphisms of moduli spaces
$\zeta:M^{0}(l,n)\to S^{[n+\frac{l+l^{2}}{2}]},\quad E\to p_{*}(E)$
$\eta:M^{0}(l,n)\to S^{[n+\frac{-l+l^{2}}{2}]},\quad E\to p_{*}(E(C)).$
The morphisms $\zeta$ and $\eta$ are actually projection morphisms of derived Grassmannians of the universal ideal sheaves over Hilbert schemes of points on $S$: let
$\mathcal{I}_{m}\in Coh(S^{[m]}\times S)$
be the universal ideal sheaf of $S^{[m]}$ and we denote
$\mathcal{I}_{m,o}:=\mathcal{I}_{m}|_{S^{[m]}\times o},\quad(\text{short for }\mathcal{I}_{o}).$
By a folklore lemma (like Proposition 2.14 of [25]), $\mathcal{I}_{o}$ is of tor amplitude $[0,1]$.
Theorem 5.6 (Theorem 4.1 of [22]).
We have a canonical isomorphism
$M^{0}(l,n)\cong Gr_{S^{[n+\frac{-l+l^{2}}{2}]}}(\mathcal{I}_{o},l),\quad M^{0}(l,n)\cong Gr_{S^{[n+\frac{l+l^{2}}{2}]}}(\mathcal{I}_{o}^{\vee}[1],l),$
where the projection morphisms are $\eta$ and $\zeta$ exactly.
6. The admissible $\mathcal{E}$-representation on perverse coherent sheaves
Given a smooth projective surface $S$ and a closed point $o\in S$ and a cohomology theory $\mathbb{H}$, we consider the grading which we call as the $l$-grading
$\mathbb{H}^{S,o}_{l}:=\bigoplus_{n\in\mathbb{Z}}\mathbb{H}^{*}(M^{0}(l,n),\mathbb{Q}),\quad\mathbb{H}^{S,o}:=\bigoplus_{l\in\mathbb{Z}}\mathbb{H}^{S,o}_{l}.$
When $l=0$, we have
$\mathbb{H}_{0}^{S,o}\cong\bigoplus_{n\in\mathbb{Z}}\mathbb{H}^{*}(S^{[n]},\mathbb{Q}).$
Other than the $l$-grading, we also have two grading: the $q$-degree by associating $\mathbb{H}^{*}(M^{0}(l,n),\mathbb{Q})$ with degree $n$, the $t$-degree (or bi-degree) which is the cohomology degree. We denote
$1_{\mathbb{H}}:=\begin{cases}0,\text{ if }\mathbb{H}\text{ is the Hochschild homology group},\\ 1,\text{ if }\mathbb{H}\text{ is the Chow group or Borel-Moore homology group},\\ (1,1)\text{ if }\mathbb{H}\text{ is the Hodge cohomology group.}\\ \text{Undefined}\text{ if }\mathbb{H}\text{ is the Grothendieck group of coherent sheaves.}\end{cases}$
In this section, we prove that
Theorem 6.1 (Theorem 1.6).
There is a $\mathcal{E}$-representation on $\mathbb{H}^{S,o}$ such that
  1. (1)
    Under the $l$-grading, the representation is graded admissible;
  2. (2)
    the $q$-degree of $E,F$ are preserved, and the $q$-degree of $K_{\pm 1}$ are ±$\pm l$ on $\mathbb{H}^{S,o}_{l}$
  3. (3)
    the $t$-degree of $E,F$ are preserved, and the $t$-degree $K_{\pm 1}$ is ±$\pm l\cdot 1_{\mathbb{H}}$ on $\mathbb{H}^{S,o}_{l}$.
  4. (4)
    the stable limit
    $\mathbb{H}_{\infty}^{S,o}\cong\bigoplus_{n\in\mathbb{Z}}\mathbb{H}^{*}(\hat{S}^{[n]},\mathbb{Q}).$
  5. (5)
    the operators are given by geometric correspondences.
Once we have the above graded admissible representation, by the categorical Boson-Fermion correspondence theorem (i.e. Theorem 2.21), we have
Theorem 6.2 (Theorem 1.3).
We have an identification
$\mathbb{H}^{S,o}\cong\mathbb{H}_{0}^{S,o}\otimes_{\mathbb{Q}}F.$
and the action of the infinite-dimensional Clifford algebra on $F$ is given by geometric correspondences. The $t$-degree of a monomial in $F$:
$x_{i_{1}}\wedge\cdots x_{i_{n}}$
is $(i_{1}+\cdots+i_{n}-\frac{n^{2}+n}{2})\cdot 1_{\mathbb{H}}$ and the $q$-degree is $i_{1}+\cdots+i_{n}-\frac{n^{2}+n}{2}$.
Theorem 6.3 (Theorem 1.1).
Let $\mathbb{H}^{*}$ be the functor of singular cohomology groups, Hochschild homology groups, or the Chow groups. We have an identification
(6.1) $\mathbb{H}_{\infty}^{S,o}\cong\mathbb{H}_{0}^{S,o}\otimes_{\mathbb{Q}}B.$
In particular, the Heisenberg algebra action on $B$ is induced by geometric correspondences. For a monomial $y_{\lambda}$ in the Bosonic Fock space, its $t$-degree is $|\lambda|\cdot 1_{\mathbb{H}}$ and the $q$-degree is $|\lambda|$.
6.1. The geometric correspondences
We recall that over $S^{[n+\frac{-l+l^{2}}{2}]}$, by Theorem 5.6 the derived Grassmannian of $\mathcal{I}_{o}$ induces
$Gr(\mathcal{I}_{o}^{\vee}[1],l)\cong M^{0}(l,n-l),\quad Gr(\mathcal{I}_{o}^{\vee}[1],l-1)\cong M^{0}(l-1,n),\quad Gr(\mathcal{I}_{o},l)\cong M^{0}(l,n).$
Thus the operators
$\psi_{l,l-1}\quad\phi_{l,l-1}\quad\psi_{l,l}\quad\phi_{l,l}$
acts on
$\psi_{l,l-1}:\bigoplus_{n\geq 0}\mathbb{H}^{*}(M^{0}(l,n),\mathbb{Q})\to\bigoplus_{n\geq 0}\mathbb{H}^{*}(M^{0}(l-1,n),\mathbb{Q}),$
$\phi_{l-1,l}:\bigoplus_{n\geq 0}\mathbb{H}^{*}(M^{0}(l-1,n),\mathbb{Q})\to\bigoplus_{n\geq 0}\mathbb{H}^{*}(M^{0}(l,n),\mathbb{Q}),$
$\psi_{l,l}:\bigoplus_{n\geq 0}\mathbb{H}^{*}(M^{0}(l,n),\mathbb{Q})\to\bigoplus_{n\geq 0}\mathbb{H}^{*}(M^{0}(l,n-l),\mathbb{Q}),$
$\phi_{l,l-1}:\bigoplus_{n\geq 0}\mathbb{H}^{*}(M^{0}(l,n-l),\mathbb{Q})\to\bigoplus_{n\geq 0}\mathbb{H}^{*}(M^{0}(l,n),\mathbb{Q}).$
We define the operators
(6.2) $E:=\bigoplus_{l\geq 0}\phi_{l-1,l},\quad F:=\bigoplus_{l\geq 0}\psi_{l-1,l},$
$K_{1}:=\bigoplus_{l\geq 0}\phi_{l,l},\quad K_{-1}:=\bigoplus_{l\geq 0}\psi_{l,l}.$
Then they all act on $\mathbb{H}_{l,n}^{S,o}$ (for the Grothendieck Grothendieck or Hochschild homology, we replace it by $\Psi_{*,*}$ and $\Phi_{*,*}$ or $[\Psi_{*,*}]$ and $[\Phi_{*,*}]$ respectively).
Proof of Theorem 6.1.
By Proposition 4.5, operators in 6.2 forms an $\mathcal{E}$-representation by geometric correspondences and the $q,t$-degree matches the Theorem 6.1. The admissibility follows from the fact that
$M(l,n)=\emptyset$
when $n<0$. ∎
Appendix A Some algebra computations
In this section, we give the proof of Proposition 2.2, Proposition A.4, Proposition 2.17 and Proposition 2.19, which involve explicit computations.
A.1. The proof of Proposition 2.2
We recall the definition of the $P_{i},Q_{i}$ for $i>0$ in the algebra $\mathcal{E}$:
Definition A.1.
We define elements $P_{i},Q_{i}$ in $\mathcal{E}$ for any $i>0$ such that $P_{1}=EK_{-1},Q_{1}=K_{1}F$ and
(A.1) $P_{i}=K_{1}P_{i-1}K_{-1}-EP_{i-1}F,\quad Q_{i}=K_{1}Q_{i-1}K_{-1}-EQ_{i-1}F,\quad\text{ if }i>1.$
Proof of Proposition 2.2.
We denote
$P(z):=\sum_{i=1}^{\infty}P_{i}z^{i},\quad Q(z):=\sum_{i=1}^{\infty}Q_{i}z^{i}.$
Then $P(z),Q(z)$ are decided by $P_{1},Q_{1}$ and the equation
(A.2) $P(z)=z(P_{1}+K_{1}P(z)K_{-1}-EP(z)F)$
(A.3) $Q(z)=z(Q_{1}+K_{1}Q(z)K_{-1}-EQ(z)F).$
We need to prove that
$P(z)P(w)+P(w)P(z)=0,$
$Q(z)Q(w)+Q(w)Q(z)=0,$
$P(z)Q(w)+Q(w)P(z)=\frac{zw}{1-zw}.$
By A.2, we have
(A.4) $P(z)K_{1}=z(E+K_{1}P(z))\quad K_{-1}P(z)=zP(z)K_{-1},$
$Q(z)K_{1}=zK_{1}Q(z),\quad K_{-1}Q(z)=z(F+K_{-1}Q(z)),$
and
(A.5) $P(z)E=-zEP(z),\quad FP(z)=z(K_{-1}-P(z)F),$
$Q(z)E=z(K_{1}-EQ(z)),\quad FQ(z)=-zQ(z)F.$
Moreover, we have
$P_{1}P(z)=EK_{-1}P(z)=zEP(z)K_{-1}=-P(z)P_{1},$
$P_{1}Q(z)=EK_{-1}Q(z)=zE(F+Q(z)K_{-1})$
$=-Q(z)P_{1}+z(EF+K_{1}K_{1})=-Q(z)P_{1}+z.$
and similarly
$Q_{1}P(z)=-P(z)Q_{1}+z,\quad Q_{1}Q(z)=-Q(z)Q_{1}.$
Hence we have
(A.6) $P(z)P(w)=zw(P_{1}+K_{1}P(z)K_{-1}-EP(z)F)(P_{1}+K_{1}P(w)K_{-1}-EP(w)F)$
$=zP_{1}P(w)+wP(z)P_{1}-zwP_{1}^{2}+zw(K_{-1}P(z)P(w)K_{1}+EP(z)P(w)F).$
We notice that $P_{1}^{2}=EK_{-1}EK_{1}=0$. Thus let $\mathfrak{P}=P(z)P(w)+P(w)P(z)$. We have
$\mathfrak{P}=z(P_{1}P(w)+P(w)P(z))+w(P_{1}P(z)+P(z)P_{1})+zw(K_{-1}\mathfrak{P}K_{1}+F\mathfrak{P}E)$
$=zw(K_{-1}\mathfrak{P}K_{1}+F\mathfrak{P}E)$
By the boundary condition, the only solution of
$\mathfrak{P}=zw(K_{-1}\mathfrak{P}K_{1}+F\mathfrak{P}E)$
is 00 and thus
$P(z)P(w)=P(w)P(z).$
Similarly, we have
$Q(z)Q(w)=Q(w)Q(z).$
Similar to A.6, we have
$P(z)Q(w)=zP_{1}Q(w)+wP(z)Q_{1}-zwP_{1}Q_{1}+zw(K_{-1}P(z)Q(w)K_{1}+EP_{-1}(z)P(w)F),$
$Q(w)P(z)=wQ_{1}P(z)+zQ(w)P_{1}-zwQ_{1}P_{1}+zw(K_{-1}Q(z)P(w)K_{1}+EQ(z)P(w)F).$
We notice that
$P_{1}Q_{1}=EF,\quad Q_{1}P_{1}=K_{-1}K$
and thus $P_{1}Q_{1}+Q_{1}P_{1}=1$. Hence let $\mathfrak{Q}:=P(z)Q(w)+Q(w)P(z)$ and we have
$\mathfrak{Q}=zw(1+K_{-1}\mathfrak{Q}K_{1}+E\mathfrak{Q}F)$
and hence
$(\mathfrak{Q}-\frac{zw}{1-zw})=zw(K_{-1}(\mathfrak{Q}-\frac{zw}{1-zw})K_{1}+E(\mathfrak{Q}-\frac{zw}{1-zw})F)$
and hence
$P(z)Q(w)+Q(w)P(z)=\frac{zw}{1-zw}.$
A.2. The $\mathcal{E}$-representation on a graded admissible $\mathcal{C}$-module
Let $W$ be a graded admissible $\mathcal{C}$-module. We formally define
$P(z):=\sum_{i=1}^{\infty}P_{i}z^{i},\quad Q(z):=\sum_{i=1}^{\infty}Q_{i}z^{i}$
For any operator $T$ on $W$, we consider
$\Gamma(T)(z):=\sum_{i\in\mathbb{Z}}\Gamma_{i}(T)z^{i}$
such that
$\Gamma_{i}(T)=\sum_{\begin{subarray}{c}m-n=i\\ m,n\geq 1\end{subarray}}P_{m}TQ_{n}$
which is well-defined due to the admissible property. We also formally written
$\Gamma(T)(z):=P(z)TQ(z^{-1}).$
We denote
$\Lambda:=\Gamma_{0}(id).$
Lemma A.2.
We have $\Lambda|_{W_{i}}=i\cdot id$, which is an invertible operator when $i>0$.
To prove Lemma A.2, we first prove that
Lemma A.3.
Let $\alpha$ and $\beta$ be two degree 00 operators on $W$. Then $\alpha=\beta$ if
$\beta|_{W_{0}}=\alpha|_{W_{0}}\quad\text{ and }\quad[Q_{l},\beta]=[Q_{l},\alpha]\text{ for all }l>0.$
Proof.
We prove that $\alpha-\beta=0$ on all $W_{i}$ by the induction on $i$. Assuming that $\alpha-\beta|_{W_{i}}=0$, then on $W_{i+1}$,
$Q_{l}(\alpha-\beta)|_{W_{i+1}}=(\alpha-\beta)Q_{l}|_{W_{i}}=0\text{ for all }l>0.$
Hence $\alpha-\beta=0$ on $W_{i+1}$ by the admissible property of $W$. ∎
Proof of Lemma A.2.
We notice that
$Q_{l}\Lambda=-\sum_{i=1}^{\infty}P_{i}Q_{l}Q_{i}+Q_{l}=\Lambda Q_{l}+Q_{l}$
and thus $\Lambda$ and $\oplus_{i\geq 0}i\cdot id$ all satisfy the equation
$[Q_{l},*]=Q_{l}$
for all $l>0$. Hence they are the same by Lemma A.3. ∎
By induction there are unique operators $K_{1}$, $K_{-1}$ on $W$ satisfy the following relations
  1. (1)
    $K_{1},K_{-1}$ are identity on $W_{0}$;
  2. (2)
    we have
    $\Lambda K_{1}=\Gamma_{1}(K_{1}),\quad\Lambda K_{-1}=\Gamma_{-1}(K_{-1}).$
We also define
$E:=P_{1}K_{1},\quad F:=K_{-1}Q_{1}.$
Proposition A.4.
The operators $K_{1},K_{-1},E,F$ induce a graded admissible $\mathcal{E}$-module on $W$.
To prove Proposition A.4, we first prove the following lemmas
Lemma A.5.
The operators $K_{1},K_{-1},E,F$ satisfy the equation A.4, i.e.
$Q_{j}K_{1}=K_{1}Q_{j-1},\quad K_{-1}P_{j}=P_{j-1}K_{-1},\text{ if }j>0$
$P_{j}K_{1}=K_{1}P_{j-1},\quad K_{-1}Q_{j}=Q_{j-1}K_{-1},\text{ if }j>1,$
where we denote $P_{0}=Q_{0}=0$.
Proof.
We only prove $Q_{j}K_{1}=K_{1}Q_{j-1}$ and other computations follow from the same method. We notice that
$(\Lambda+id)Q_{j}K_{1}=Q_{j}\Lambda K_{1}=\sum_{l=1}^{\infty}Q_{j}P_{l+1}K_{1}Q_{l}=-\sum_{l=1}^{\infty}P_{l+1}(Q_{j}K_{1})Q_{l}+K_{1}Q_{j-1}$
$\Lambda K_{1}Q_{j-1}=\sum_{l=1}^{\infty}P_{l+1}K_{1}Q_{l}Q_{j-1}=-\sum_{l=1}^{\infty}P_{l+1}(K_{1}Q_{j-1})Q_{l}$
Thus $\alpha:=Q_{j}K_{1}-K_{1}Q_{j-1}$ satisfy the equation that
(A.7) $(\Lambda+id)\alpha=-\Gamma_{1}(\alpha).$
By induction, the solutions of A.7 is decided by $\alpha|_{W_{0}}$, and hence $\alpha=0$ follows from $\alpha|_{W_{0}}=0$. ∎
Corollary A.6.
For any $j>0$ (resp. $j>1$), we have
$Q_{j}K_{-1}K_{1}=K_{-1}K_{1}Q_{j},\quad(\text{resp. }Q_{j}K_{1}K_{-1}=K_{1}K_{-1}Q_{j}),$
i.e.
(A.8) $[Q(z),K_{-1}K_{1}]=0,\quad[\frac{d}{dz}Q(z),K_{1}K_{-1}]=0.$
Proof.
It directly follows from Lemma A.5. ∎
Corollary A.7.
On $W$ we have
$K_{1}K_{-1}=Q_{1}P_{1},\quad K_{-1}K_{1}=id.$
Proof.
The second formula follows from A.8 and Lemma A.3. For the first formula, let $\beta:=K_{1}K_{-1}-Q_{1}P_{1}$. We have
$\beta|_{W_{0}}=0,\quad Q_{1}\beta=0,\text{ and }\quad[Q_{l},\beta]=0\text{ for all }l>1.$
Then by induction and the similar method with Lemma A.3, we also have $\beta=0$. ∎
Proof of Proposition A.4.
The formula $K_{-1}K_{1}=id$ follows from Corollary A.7. We have
$FE=K_{-1}P_{1}Q_{1}K_{1}=K_{-1}K_{1}K_{-1}K_{1}=id$
by Corollary A.7. Also by Corollary A.7,
$K_{1}K_{-1}+EF=Q_{1}P_{1}+P_{1}Q_{1}P_{1}Q_{1}=Q_{1}P_{1}+P_{1}Q_{1}-P_{1}^{2}Q_{1}^{2}=id.$
Finally, we have
$FK_{1}=K_{-1}P_{1}K_{1}=0,\quad K_{-1}E=K_{-1}Q_{1}K_{1}=0$
by Lemma A.5. ∎
A.3. The proof of Proposition 2.17
Let $H_{\pm k}$, $k\in\mathbb{Z}-\{0\}$ be the generators of the infinite-dimensional Heisenberg algebra.
Let $W_{\infty}$ be an admissible $\mathcal{H}$-representation, with the filtration
$W_{i}:=\bigcap_{l>i}ker(H_{l}).$
and the operators for every $k>0$
$\mathcal{L}_{k}:=-\sum_{i=1}^{\infty}\frac{1}{k^{i}i!}H_{k}^{i-1}H_{-k}^{i}.$
Lemma A.8.
We have
(A.9) $\mathcal{L}_{k}H_{k}=id,\quad H_{-k}(1-H_{k}\mathcal{L}_{k})=0.$
Proof.
By induction, for any integer $i\geq 0$, we have
(A.10) $H_{-k}H_{k}^{i}=H_{k}^{i}H_{-k}-ik,\quad H_{-k}^{i}H_{k}=H_{k}H_{-k}^{i}-ik.$
By A.10 we have
$\mathcal{L}_{k}H_{k}=-\sum_{i\geq 1}\frac{1}{k^{i}i!}H_{k}^{i-1}H_{-k}^{i}H_{k}=\sum_{i\geq 1}(\frac{1}{k^{i-1}(i-1)!}H_{k}^{i-1}H_{-k}^{i-1}-\frac{1}{k^{i}i!}H_{k}^{i}H_{-k}^{i})=1.$
and
$H_{-k}(1-H_{k}\mathcal{L}_{k})=H_{-k}(Id+\sum_{i=1}^{\infty}\frac{1}{k^{i}i!}H_{k}^{i}H_{-k}^{i})$
$=H_{-k}+\sum_{i=1}^{\infty}(\frac{1}{k^{i}i!}H_{k}^{i}H_{-k}^{i+1}-\frac{1}{k^{i-1}(i-1)!}H_{k}^{i-1}H_{-k}^{i})=H_{-k}-H_{-k}=0.$
Proof of Proposition 2.17.
We only need to prove that it forms a graded $\mathcal{E}$-representation, and the admissibility follows from the admissibility of $\mathcal{W}_{\infty}$.
First, we notice that for any $k>0$, the restriction of
$\mathcal{L}_{k}:W_{k}\to W_{k}$
at $W_{k-1}$ is zero. Thus we have
$K_{-1}E=0,\quad FE=id.$
The formula $K_{1}K_{-1}+EF=0$ follows from the definition of $K_{\pm 1}$ and $E,F$.
On the other hand, by A.9 we have
$K_{-1}K_{1}=id,\quad FK_{1}=(1-K_{1}K_{-1})K_{1}=0.$
Thus $W$ is a graded $\mathcal{E}$-representation. ∎
A.4. The proof of Proposition 2.19
Let $W$ be an admissible graded $\mathcal{E}$-representation. As the limit of 2.2, on $W_{l}$ ($l>0$) we have
(A.11) $\sum_{l=0}^{\infty}E_{l}F_{l}=id.$
Lemma A.9.
For any $i>0$, the operators $H_{i,i}$ and $H_{i,-i}$ satisfy the relation
$H_{i,i}H_{i-i}-H_{i,-i}H_{i,i}=i\cdot id.$
Proof.
We have
$H_{i,i}H_{i,-i}=-iK_{1}\sum_{l=0}^{\infty}K_{1}^{l}K_{-1}^{l+1}=i-\sum_{l=0}^{\infty}(K_{1})^{l}K_{-1}^{l}$
$=i-\sum_{l=0}^{\infty}(K_{1})^{l}K_{-1}^{l+1}K_{1}=i+H_{i,-i}H_{i,i}.$
Lemma A.10.
For any $i>0$ and $|j|<i$, we have
$H_{i,j}H_{i,i}=H_{i,i}H_{i,j},\quad H_{i,j}H_{i,-i}=H_{i,-i}H_{i,j}$
Proof.
First we prove that $H_{i,j}H_{i,i}=H_{i,i}H_{i,j}$. It follows from
$H_{i,j}H_{i,i}=\sum_{l=0}^{\infty}E_{l}H_{i-1,j}F_{l}K_{1}=\sum_{l=0}^{\infty}E_{l+1}H_{i-1,j}F_{l}=H_{i,i}H_{i,j}$
by Lemma 2.1. Then we prove that $H_{i,j}H_{i,-i}=H_{i,-i}H_{i,j}$. It follows from
$H_{i,j}H_{i,-i}=-i\sum_{l,m\geq 0}E_{l}H_{i-1,j}F_{l}K_{1}^{m}K_{-1}^{m+1}=-i\sum_{l\geq m\geq 0}E_{l}H_{i-1,j}F_{l+1},$
$H_{i,-i}H_{i,j}=-i\sum_{l,m\geq 0}K_{1}^{m}K_{-1}^{m+1}E_{l}H_{i-1,j}F_{l}=-i\sum_{l>m\geq 0}E_{l-1}H_{i-1,j}F_{l}$
$=-i\sum_{l\geq m\geq 0}E_{l}H_{i-1,j}F_{l+1}=H_{i,j}H_{i,-i}$
by Lemma 2.1. ∎
Proposition A.11.
On $W_{i}$, the operators $H_{i,j}$ where $j\neq 0$ and $-i\leq j\leq i$ satisfy the relation
(A.12) $H_{i,j}H_{i,k}-H_{i,k}H_{i,j}=j\delta_{j,-k}id.$
Proof.
We assume that $|j|\leq|k|$. We prove it by induction on $i$. It is trivial when $i=0$. Assuming it is true for $i-1$, then if $k=\pm i$, A.12 follows from Lemma A.9 and Lemma A.10. Otherwise, by Lemma 2.1 we have
$H_{i,j}H_{i,k}=\sum_{l,m\geq 0}E_{l}H_{i-1,j}F_{l}E_{m}H_{i-1,k}F_{m}=E_{l}H_{i-1,j}H_{i-1,k}F_{l}.$
Thus we have
$[H_{i,j},H_{i,k}]=\sum_{l=0}^{\infty}E_{l}[H_{i-1,j},H_{i-1,k}]F_{l}=j\delta_{j,-k}\sum_{l=0}^{\infty}E_{l}F_{l}=j\delta_{j,-k}id$
by A.11. ∎
Proof of Proposition 2.19 .
It is the limit of Proposition A.11 when $i\to\infty$. ∎